### v3

# FusionBench: A Comprehensive Benchmark of Deep Model Fusion

Abstract
Deep model fusion is an emerging technique that unifies the predictions or parameters of several deep neural networks into a single model in a cost-effective and data-efficient manner. This enables the unified model to take advantage of the original models’ strengths, potentially exceeding their performance. Although a variety of deep model fusion techniques have been introduced, their evaluations tend to be inconsistent and often inadequate to validate their effectiveness and robustness against distribution shifts. To address this issue, we introduce FusionBench, which is the first comprehensive benchmark dedicated to deep model fusion. FusionBench covers a wide range of tasks, including open-vocabulary image classification, text classification, and text-to-text generation. Each category includes up to eight tasks with corresponding task-specific models, featuring both full fine-tuning and LoRA fine-tuning, as well as models of different sizes, to ensure fair and balanced comparisons of various multi-task model fusion techniques across different tasks, model scales, and fine-tuning strategies. We implement and evaluate a broad spectrum of deep model fusion techniques. These techniques range from model ensemble methods, which combine the predictions to improve the overall performance, to model merging, which integrates different models into a single one, and model mixing methods, which upscale or recombine the components of the original models. FusionBench now contains 26 distinct tasks, 74 fine-tuned models, and 16 fusion techniques, and we are committed to consistently expanding the benchmark with more tasks, models, and fusion techniques. In addition, we offer a well-documented set of resources and guidelines to aid researchers in understanding and replicating the benchmark results. This includes detailed documentation, code examples, and tutorials, making FusionBench a user-friendly and accessible platform for both beginners and experienced researchers. The FusionBench homepage is available at https://github.com/tanganke/fusion_bench.
1 Introduction
In recent years, a new paradigm called “learn from model” has emerged in the field of deep learning, which focuses on leveraging the knowledge embedded in existing models to develop new ones [67]. This paradigm has been widely adopted in various scenarios, such as model tuning [18, 8], model distillation [21], model pruning [16, 3], model editing [38, 66], and so on. Among these methods, deep model fusion is particularly appealing. It merges the parameters or predictions of multiple models to create a more robust and efficient unified model. Due to its effectiveness and scalability, many new techniques for deep model fusion have recently been proposed [34].
Deep model fusion offers both scalability and data efficiency by utilizing the knowledge embedded in pre-existing models, rather than requiring training from scratch. This approach significantly accelerates model development, making it a practical solution in the current era dominated by large foundation models. Despite its potential, the evaluation of deep model fusion techniques often suffers from inconsistency and inadequacy. Standardized assessments are lacking, making it challenging to verify their effectiveness and robustness. The potential reasons for this inconsistency include the rapid development of new techniques, the absence of standardized tasks and models, and the variety of settings (such as different fine-tuning strategies). Additionally, challenges in implementing or replicating prior work contribute to these inconsistencies.
To tackle these challenges, we develop the first comprehensive benchmark dedicated to deep model fusion, called FusionBench. FusionBench is built to be a modular and extensible platform, comprising three core modules: the Algorithm Module, the Model Pool Module, and the Task Pool Module. Each module is configurable, allowing users to easily customize and manage the benchmark settings. In summary, our main contributions are four-fold:
  1. 1.
    An extensive codebase and configurable interface: Our codebase is structured around three core modules: the algorithm module, the model pool module, and the task pool module. Moreover, to facilitate ease of use and flexibility, we provide a unified command line interface with YAML configuration options for effortless customization.
  2. 2.
    Comprehensive evaluations: We conduct thorough evaluations of the deep model fusion techniques, assessing their performance across a variety of benchmarks and settings.
  3. 3.
    Extensive analysis and insights: We provide detailed analysis and insights based on the evaluation results, revealing key factors that influence their effectiveness. This includes identifying best practices, exploring the impact of fine-tuning strategies, and offering recommendations for future research.
  4. 4.
    User-friendly resources and guidelines: We offer a well-documented set of resources and guidelines to aid researchers in understanding and replicating the benchmark results. This includes detailed documentation, code examples, and tutorials, making FusionBench a user-friendly and accessible platform for both beginners and experienced researchers. Documentation is available at https://tanganke.github.io/fusion_bench/.
2 Related Work
Figure 1: A taxonomy of deep model fusion techniques.
Since deep model fusion is a relatively new research area, there is currently no standardized taxonomy. Different researchers may categorize these techniques in various ways based on their understanding and points of view. Here, we propose a taxonomy that divides these techniques into three major categories: Model Ensemble, Model Merging, and Model Mixing. Each of these categories approaches model fusion from a unique perspective, offering distinct advantages and applicability. In the following, we provide detailed explanations, formal definitions, and analyze their strengths and weaknesses. A visualization of the taxonomy is shown in Figure 1.
Model Ensemble methods combine the predictions of multiple models to improve the overall performance of a machine learning system [45], where the collective knowledge of multiple models is often more accurate and reliable than that of any individual model. Mathematically, given a set of $N$ models $\{f_{1},f_{2},\dots,f_{N}\}$, which can be homogeneous or heterogeneous, we use their predictions to obtain a global prediction $y=\mathcal{A}(x;f_{1},f_{2},\dots,f_{N};w)$, where $\mathcal{A}$ is an ensemble algorithm and $w$ are the algorithmic parameters. Each model $f_{i}$ can also be associated with a specification to indicate its weight or importance in the ensemble [41, 68, 60, 49]. Ensemble methods are widely used and effective in improving performance but are often expensive to use and manage. Recently, there have also been studies exploring efficient methods for model ensembles [55, 5, 2].
Model Merging methods integrate the parameters of multiple models into a single model, which can be more efficient and scalable model fusion techniques. Mathematically, given a set of $N$ isomorphic models $\{f_{i}(\cdot;\theta_{i})\}_{i=1}^{N}$, each parameterized with $\theta_{i}$, we merge them into a single model with parameters $\theta=\mathcal{A}(\theta_{1},\theta_{2},\dots,\theta_{N};w)$, where $\mathcal{A}$ is a merging algorithm and $w$ are the algorithmic parameters. The merged model can be expressed as $f(\cdot;\theta)$. This method can be implemented through linear interpolation in parameter space [57, 25, 62, 37, 65, 7, 44, 15, 36], leveraging mode connectivity [10, 11, 4, 12, 42], aligning features or parameters [35, 1, 28, 48, 13, 26], and ensemble distillation [52, 53]. Model merging methods are often performed in a data-efficient manner, the algorithmic parameters $w$ can also be learned during test time via test-time adaptation training or meta-learning for a more seamless merging [63, 50].
Model Mixing methods fuse the components of multiple models to create a new model, which can be more flexible and adaptive than the original models. Mathematically, given a set of $N$ models $\{f_{i}(\cdot;\theta_{i})\}_{i=1}^{N}$, each parameterized with $\theta_{i}$, we mix their components to obtain a new model with parameters $\Theta=\mathcal{A}(\theta_{1},\theta_{2},\dots,\theta_{N};w)$, where $\mathcal{A}$ is a mixing algorithm and $w$ is the algorithmic parameters. The mixed model can be expressed as $F(\cdot;\Theta)$, which often has more parameters than the original models, and thus can be more expressive and powerful to capture the underlying patterns in the data. Model mixing methods can be implemented through layer recombinations [23, 27], model stitching [33, 39], or upscale to create a Mixture of Experts (MoE)-based sparse model [30, 64, 51].
Although several model fusion methods have been proposed, benchmarks and unified toolkits are still lacking in this field. A recent notable work, MergeKit [14], provides a collection of model fusion techniques specifically designed for merging large language models (LLMs), with a focus on model merging methods and Transformer-based LLMs. However, MergeKit’s scope is limited to a specific domain and model architecture. In contrast, FusionBench aims to be a comprehensive benchmark that encompasses a wide range of tasks and model types. It includes a diverse set of fine-tuned models and tasks to evaluate, making it a more generalized and versatile platform for assessing the performance of different model fusion approaches across various domains and architectures.
3 Our Benchmark
Figure 2: The general framework of the modularized FusionBench codebase.
The general framework of the modularized FusionBench codebase is shown in Figure 2, which consists of three primary elements: Algorithm Module, Model Pool Module, and Task Pool Module. In Section 3.1, we introduce the codebase, which is designed to be flexible and modular, allowing users to easily run experiments and evaluate the performance of model fusion algorithms. In Section 3.2 and Section 3.3, we introduce the implemented model fusion algorithms and the tasks and models included in FusionBench. Finally, in Section 3.4, we discuss the documentation and tutorials provided to help users understand the benchmark and effectively use the codebase.
3.1 Codebase
We’ve constructed a flexible and modular codebase, which serves as the foundation for FusionBench. As shown in Figure 2, the codebase is composed of three primary elements: Algorithm Module, Model Pool Module, and Task Pool Module, which are responsible for implementing the model fusion algorithms, managing the models to be fused, and managing the tasks to be evaluated, respectively. Additionally, we provide a command line interface (CLI) to facilitate the use of the codebase and to enable users to easily run experiments and evaluate the performance of model fusion algorithms.
Algorithm Module is the core component of the codebase, which contains the implementation of various model fusion algorithms. Each algorithm is implemented as a separate Python class, which inherits from the base class ModelFusionAlgorithm. The algorithm classes are designed to be configurable and independently callable, allowing users to easily instantiate and set up the algorithms through our CLI or by directly invoking the Python classes in their own code.
Model Pool Module is responsible for managing the models to be fused. It offers a unified interface for loading the pre-trained model and fine-tuned models. The module is designed to be extensible, allowing users to easily add support for new model architectures and add their own models to the pool. Each model in the pool can also be associated with metadata to meet the requirements of specific model fusion algorithms, such as the test dataset for test-time adaptation training.
Task Pool Module is responsible for managing the tasks to be evaluated. Each task comprises a dataset and a set of evaluation metrics, which are defined in the YAML configuration file. This module offers a unified interface for loading tasks and assessing the performance of model fusion algorithms on these tasks. Users can effortlessly add support for new task types and evaluation metrics, or add new tasks of the same type but with different datasets.
3.2 Implemented Algorithms
Table 1: Implemented model fusion algorithms in FusionBench.
Category Method Requirement
Model Ensemble Simple Ensemble [45] -
Weighted Ensemble [45] hyperparameter search
Max-Model Predictor [59] -
Model Merging Simple Average / Modelsoups [57] -
Weighted Average [37] hyperparameter search
Fisher Merging [37] compute weights on labeled data
RegMean [28] compute weights on labeled data
Task Arithmetic [25] hyperparameter search
Ties-Merging [62] hyperparameter search
Task-Wise AdaMerging [63] test-time adaptation training
Layer-Wise AdaMerging [63] test-time adaptation training
Model Mixing Depth Upscaling [29] pre-training to recover performance
MoE-based Upscaling [30] pre-training to recover performance
MoE-based Merging [30] training on the combined model
Weight-Ensemble MoE [51] test-time adaptation training, vision tasks
Model Recombination [23] training on the combined model
In our benchmark, we have implemented 16 model fusion algorithms as the initial set. This includes 3 model ensemble methods, 8 model merging methods, and 5 model mixing methods. Our primary selection criterion for choosing among various algorithms is their applicability and effectiveness within the realm of deep learning architectures We have also considered the popularity of the algorithms in the literature and their practical applicability, such as their potential use in large-scale language models. We list the implemented algorithms in Table 1.
As shown in Table 1, we implemented three kinds of model fusion algorithms. A brief introduction and formal definition of our taxonomy are provided in Section 2. Model ensemble methods are effective at enhancing the performance of a machine learning system, but they are computationally expensive to infer. Model merging methods aim to integrate the advantages of individual models, making them popular in multi-task model fusion and auxiliary learning. In these scenarios, multiple single-task models are merged to construct a multi-task model, or models focused on auxiliary tasks are combined to boost the performance of a primary task. Model mixing methods are frequently used to scale up a pre-trained model to a larger size or to combine multiple models into a new one. Consequently, model mixing methods often necessitate additional training after the fusion process.
3.3 Tasks and Models
Table 2: Tasks and models included in FusionBench for evaluating multi-task model fusion algorithms.
Domain Task Type Datasets Models
Computer Vision Image classification (8 tasks) SUN397 [61], Stanford Cars [31], RESISC45 [6], EuroSAT [19], SVHN [40], GTSRB [47], MNIST [32], DTD [9] 8$\times$×CLIP-ViT-B/32 (Transformers [56]), 8$\times$×CLIP-ViT-B/32 (OpenCLIP [24]), 8$\times$×CLIP-ViT-B/16 (OpenCLIP [24]), 8$\times$×CLIP-ViT-L/14 (Transformers [56]), 8$\times$×CLIP-ViT-L/14 (OpenCLIP [24])
Sence Understanding (3 tasks) NYUv2 [46] 3$\times$×Resnet-50 models
Natural Language Processing Text classification (7 tasks) CoLA, MNLI, MRPC, QNLI, QQP, RTE, and SST-2 [54] 7$\times$×GPT-2
Text-to-text generation (8 tasks) CoLA, MNLI, MRPC, QNLI, QQP, RTE, SST-2, and STSB [54] 16$\times$×Flan-T5-Base (w/ & w/o LoRA), 8$\times$×Flan-T5-Large (w/ LoRA)
(a) Clean
(b) Motion
(c) Impulse
(d) Gaussian
(e) Pixelate
(f) Spatter
(g) Contrast
(h) JPEG
Figure 3: Here are eight instances of distorted images from the Stanford Cars dataset. We use these distortions to assess the robustness and generalization capacity of the merged model as in [63, 51].
Model fusion is a versatile technique that can be applied across various machine learning tasks at different stages of model development. In FusionBench, we specifically provide a diverse array of tasks and corresponding fine-tuned models to ensure a fair and comprehensive evaluation of multi-task model fusion algorithms. We have selected tasks from the domains of computer vision and natural language processing, as these are the most popular and extensively studied areas in deep learning research. The tasks included in our benchmark are open-vocabulary image classification, text classification, and text-to-text generation. We list these tasks and models in Table 2. We make them publicly available to facilitate reproducibility and further research at HuggingFace  [^1] Access to datasets and fine-tuned models at https://huggingface.co/tanganke.
Open-vocabulary image classification is a challenging task that requires models to classify images into a large number of categories. We have selected eight image classification datasets, including SUN397 [61], Stanford Cars [31], RESISC45 [6], EuroSAT [19], SVHN [40], GTSRB [47], MNIST [32], DTD [9]. These datasets cover a wide range of image classification tasks, including object recognition, satellite image classification, and texture classification. We fine-tuned two CLIP-ViT models, CLIP-ViT-B/32 and CLIP-ViT-L/14, on these datasets. We report accuracy as the evaluation metric for these tasks. Specifically, to assess the robustness of multi-task model fusion algorithms, particularly those needing test-time adaptation training, we adopt the techniques recommended by Hendrycks and Dietterich [20] to create corrupted versions of the test set for Cars, EuroSAT, RESISC45, and GTSRB. These corruptions are designed to simulate common image corruptions in real-world scenarios, including motion blur, impulse noise, Gaussian noise, pixelation, spatter, contrast adjustments, and JPEG compression.
Scene understanding tasks are performed using the NYUv2 [46] dataset, which consists of RGB-D images and includes three tasks: 13-class segmentation, depth estimation, and surface normal estimation. We fine-tuned ResNet-50 models [17] as the backbone for our experiments. The initial weights for these models were pre-trained on the ImageNet dataset. We then adapted them to the specific tasks of segmentation, depth estimation, and surface normal estimation.
Text classification is a fundamental task in natural language processing that involves categorizing text data into predefined classes. We have selected seven text classification tasks from the General Language Understanding Evaluation (GLUE) benchmark [54], including CoLA, MNLI, MRPC, QNLI, QQP, RTE, and SST-2. We fine-tuned GPT-2 models on these seven tasks, each with a different head for classification [43]. We report accuracy as the evaluation metric.
Text-to-text generation is more challenging than text classification, as it requires models to generate appropriate text outputs instead of projecting the hidden representations to logits. Similar to text classification, we have selected eight text-to-text generation tasks from the GLUE benchmark, including CoLA, MNLI, MRPC, QNLI, QQP, RTE, SST-2, and STSB. We fine-tuned Flan-T5 models on these tasks, with and without the LoRA adaptation [22]. The prompt templates for these tasks are provided in Appendix D. We report Spearman’s $\rho$ for STSB and exact match accuracy for other tasks. In particular for STSB, if text outputs can’t be parsed as numerical values, we report the $\rho$ as 0.
3.4 Documentation and Tutorials
Documentation and tutorials are essential for beginners to understand the methodology behind the benchmark, to reproduce the experiments, and to effectively use the codebase. To this end, we offer comprehensive documentation and tutorials on the project homepage, which guide users through the fundamentals of model fusion, the steps to run experiments, and the procedures for evaluating the performance of model fusion algorithms. Additionally, we present some experimental results to shed light on the performance of different model fusion algorithms across various tasks.
4 Evaluation and Analysis
In this section, we evaluate the performance of multi-task model fusion algorithms on a variety of tasks, as well as analyze the generalization and robustness of these algorithms. We also provide an ablation study to investigate the impact of hyperparameter selection. Most of the experiments are conducted with a single NVIDIA RTX 3090 GPU with 24GB memory.
4.1 Experimental Setup
In this section, we conduct a series of multi-task model fusion experiments on image classification tasks, text classification tasks, and text-to-text generation tasks to evaluate the performance of multi-task model fusion algorithms. These tasks are chosen to cover a wide range of NLP and CV tasks, as described in Section 3.3. Table 2 provides a summary of the tasks and models used in our experiments.
4.2 Multi-Task Model Fusion
Figure 4: Illustration of multi-task model fusion.
Table 3: Multi-task performance when merging CLIP-ViT-B/32 models on all eight tasks.
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
Reference Methods
Pre-trained 63.2 59.8 60.7 46.0 31.6 32.5 48.2 43.9 48.2
Fine-tuned (STL) 75.0 78.3 95.2 99.0 97.3 98.9 99.6 79.7 90.3
Multi-Task Model Fusion Methods
Weight Averaging [57] 65.4 62.6 70.8 76.9 64.5 54.9 86.3 50.9 66.5
Fisher Merging [37] 66.7 64.0 72.2 91.6 69.0 64.3 83.5 53.7 70.6
RegMean [28] 67.8 68.9 82.5 94.4 90.6 79.2 97.6 63.2 80.5
Task Arithmetic [25] 57.1 55.7 64.9 76.7 77.9 68.5 96.1 47.2 68.0
Ties-Merging [62] 67.1 64.2 74.1 76.8 77.7 69.4 94.1 54.0 72.2
task-wise AdaMerging [63] 58.6 56.9 69.8 82.4 70.3 58.9 97.2 55.3 68.7
layer-wise AdaMerging [63] 67.9 71.3 83.5 92.7 87.4 92.9 98.2 67.0 82.6
WEMoE (Model Mixing) [51] 73.7 76.8 93.4 98.2 96.8 98.2 99.6 76.6 89.2
Table 4: Experimental results of merging single-task Resnet50 models on three NYUv2 tasks.
Method Segmentation Depth Estimation Normal
mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Mean $\downarrow$
Single-Task Learning
Segmentation 52.0 73.8 242.8 88.7 82.8
Depth Estimation 2.3 6.2 42.5 17.7 82.8
Normal 2.0 4.9 264.0 98.1 24.7
Multi-Task Model Fusion Methods
Weight Averaging [57] 39.0 67.0 55.1 22.7 30.4
Task Arithmetic ($\lambda=0.3$[25] 33.6 63.3 56.3 23.2 31.3
Ties-Merging ($\lambda=0.3$[62] 36.3 61.7 60.5 24.5 33.1
Table 5: Generalization results on two unseen tasks when merging ViT-B/32 models on six tasks.
Method Seen Tasks Unseen Tasks
SUN397 Cars RESISC45 DTD SVHN GTSRB Avg. MNIST EuroSAT Avg.
Pre-trained 63.2 59.9 60.6 43.9 23.5 30.4 46.9 47.6 45.6 46.6
Fisher Merging [37] 65.5 67.2 78.2 57.6 84.2 75.9 71.4 71.8 49.4 60.6
RegMean [28] 68.7 70.0 86.5 65.9 93.9 86.7 78.6 82.2 49.3 65.7
Task Arithmetic [25] 64.3 63.0 73.2 54.9 84.7 79.5 69.9 75.5 42.6 59.1
Ties-Merging [62] 68.3 65.5 76.9 54.9 75.4 72.0 68.9 73.1 47.3 60.2
AdaMerging [63] 68.4 71.9 87.9 69.1 92.2 93.8 80.5 77.7 47.3 62.5
WEMoE [51] 75.4 77.5 94.3 77.0 96.8 98.7 86.6 78.3 44.0 61.1
Table 6: Ablations of the test data distribution on ViT-B/32 (for all methods, $\lambda=0.3$).
Method Cars EuroSAT RESISC45 GTSRB Avg. Cars EuroSAT RESISC45 GTSRB Avg.
Clean Test Set Corrupted Test Set (Motion Blur)
Fisher Merging [37] 66.0 92.7 83.7 78.7 80.3 60.7 57.6 81.7 78.4 69.6
RegMean [28] 72.1 97.5 88.9 93.9 88.1 70.0 71.3 87.5 86.8 78.9
Task Arithmetic [25] 64.6 91.8 80.2 74.8 77.9 62.4 59.2 78.5 63.3 65.9
Ties-Merging [62] 65.2 83.3 78.1 67.4 73.5 64.4 53.9 76.4 57.1 62.9
AdaMerging [63] 75.2 94.3 87.6 96.7 88.5 72.4 72.7 85.3 94.3 81.2
WEMoE [51] 77.4 98.9 94.4 99.0 92.4 76.5 74.2 93.7 97.4 85.5
Corrupted Test Set (Impluse Noise) Corrupted Test Set (Gaussian Noise)
Fisher Merging [37] 61.5 50.0 74.7 52.6 59.7 61.6 48.1 76.0 51.3 59.3
RegMean [28] 66.9 51.0 80.6 68.7 66.8 69.4 41.8 84.0 67.7 65.7
Task Arithmetic [25] 59.8 53.3 72.3 45.0 57.6 61.5 52.5 75.0 50.1 59.8
Ties-Merging [62] 60.2 45.6 69.8 38.3 53.5 61.8 47.3 73.1 42.3 56.1
AdaMerging [63] 69.2 40.0 79.6 83.3 68.0 70.0 53.3 82.1 80.0 71.4
WEMoE [51] 75.1 9.7 91.5 91.8 67.0 76.5 9.6 92.7 88.7 66.8
Corrupted Test Set (Pixelate) Corrupted Test Set (Spatter)
Fisher Merging [37] 2.2 34.0 17.0 63.2 29.1 61.4 64.2 74.6 47.3 61.9
RegMean [28] 2.3 38.3 18.2 89.4 37.0 67.7 60.0 81.3 81.9 72.7
Task Arithmetic [25] 2.3 33.2 19.1 65.6 30.0 61.0 62.5 72.8 57.0 63.3
Ties-Merging [62] 3.3 31.8 18.0 58.5 27.9 61.3 52.9 70.3 48.1 58.2
AdaMerging [63] 1.3 52.9 21.0 91.0 41.5 68.4 55.9 78.3 92.3 73.7
WEMoE [51] 0.5 11.6 2.3 97.5 28.0 75.1 9.7 91.4 96.3 68.1
Corrupted Test Set (Contrast) Corrupted Test Set (JPEG Compression)
Fisher Merging [37] 63.8 58.4 75.5 70.4 67.0 66.3 67.6 82.6 58.9 68.8
RegMean [28] 69.6 64.8 84.4 90.0 77.2 71.5 72.6 88.7 82.2 78.7
Task Arithmetic [25] 62.3 55.7 75.3 70.8 66.0 63.9 66.1 80.1 61.0 67.8
Ties-Merging [62] 64.2 52.4 74.8 63.5 63.7 65.0 59.5 77.9 53.2 63.9
AdaMerging [63] 73.1 67.4 83.0 96.2 79.9 72.9 70.7 86.3 90.6 80.1
WEMoE [51] 77.2 34.7 93.1 98.4 75.9 77.3 61.0 94.1 95.7 82.0
In this evaluation, we begin by comparing multi-task model fusion algorithms under several settings:
  1. 1.
    Image Classification: We use CLIP models from the HuggingFace Library [24]. The results are in Table 3 and Figure 4 for CLIP-ViT-B/32, and in Table 14 for CLIP-ViT-L/14.
  2. 2.
    Scene Understanding: Using ResNet-50 models on the NYUv2 dataset for segmentation, depth estimation, and normal estimation tasks. Results are in Table 4.
  3. 3.
    Text Classification: Results for GPT-2 models on seven tasks are shown in Table 12.
  4. 4.
    Text-to-Text Generation: For LoRA fine-tuned Flan-T5-base and Flan-T5-large models, after merging and unloading the LoRA adapters, results are in Tables 13 and 15.
In these tables, we compare the performance of different multi-task model fusion algorithms across various tasks. Pre-trained models’ performance, fine-tuned models’ performance, and traditional multi-task learning (MTL) methods are provided for reference. In Appendix A, we provide a detailed description of these fine-tuned single-task models.
We have the following key observations: (1) Multi-task model fusion algorithms generally outperform pre-trained models. This indicates that multi-task model fusion can effectively transfer knowledge from multiple single-task models to improve performance on each task. (2) Layer-wise AdaMerging and Weight-Ensembling MoE achieve the best overall performance among the multi-task model fusion methods. These methods demonstrate the effectiveness of adaptive layer-wise merging and mixture-of-experts approaches in multi-task model fusion. (3) The performance gap between multi-task model fusion and fine-tuned models (STL) is larger for CLIP-ViT-B/32 compared to CLIP-ViT-L/14. This suggests that multi-task model fusion may be more beneficial for smaller models, as they have more room for improvement through knowledge transfer. (4) Traditional MTL outperforms most multi-task model fusion methods, which indicates that traditional MTL is still a strong baseline for multi-task learning, and there is room for improvement in multi-task model fusion algorithms.
4.3 Generalization and Robustness Evaluation
To further assess the generalization and robustness of multi-task model fusion algorithms, we conduct experiments on unseen tasks and corrupted test sets (or out-of-distribution test sets). (1) Tables 5 and 16 present the generalization performance of various multi-task model fusion algorithms when merging CLIP-ViT-B/32 models trained on six seen tasks and evaluating their performance on two unseen tasks. This analysis helps us understand how well the fused models can adapt to new tasks that were not encountered during the training and model fusion process. Additional details and discussions regarding the generalization experiments can be found in Appendix C. (2) Furthermore, in Table 6, we investigate the robustness of multi-task model fusion algorithms by evaluating their performance on corrupted test sets. These corrupted test sets are designed to simulate real-world scenarios where the input data may be noisy or corrupted.
We have the following key observations: (1) The performance of all multi-task model fusion methods on unseen tasks is generally lower than their performance on seen tasks. This is expected, as the models being fused are not explicitly trained on the unseen tasks. (2) There is a negative transfer observed on the RESISC45 dataset in the generalization experiments. The performance of all multi-task model fusion methods on RESISC45 is lower than the pre-trained model, indicating that the knowledge transferred from the seen tasks may not be beneficial or even harmful to this specific unseen task. (3) The performance of all methods drops significantly on certain types of corruptions, such as pixelation and impulse noise. This highlights the challenge of maintaining robustness under severe distribution shifts and the need for further research in this direction. (4) When the test distribution is corrupted, adaptive methods may overfit to certain tasks, leading to a decrease in overall performance. This suggests that adaptive methods may need to be further regularized to improve generalization and robustness.
5 Conclusions, Future Plans and Societal Impacts
Conclusions. We’ve developed a flexible and modular codebase, which serves as the foundation for FusionBench. Our benchmark provides a comprehensive evaluation framework for assessing the performance of multi-task model fusion algorithms. This innovative and comprehensive framework underscores the advantages of a scalable and extendable architecture, thereby simplifying the creation of deep model fusion algorithms. We also organize and provide a collection of datasets and models, which can be utilized to ensure a fair comparison. Last, FusionBench comes with extensive documentation and a series of tutorials, making it user-friendly for beginners and interested researchers. We hope that the community will leverage this benchmark to develop and evaluate new fusion algorithms and to further the popularity of deep model fusion in the machine learning community.
Limitations and Future Plans. To date, FusionBench primarily focuses on the evaluation of deep model fusion algorithms for multi-task learning. Despite having implemented numerous fusion algorithms, including those that don’t primarily focus on multi-task learning, we have not yet to investigate the evaluation for these methods. In the future, we plan to extend the benchmark to provide a more comprehensive evaluation framework for them. What’s more, we plan to extend the benchmark by incorporating additional datasets and applications, such as human preference alignment, multi-modal fusion, and reinforcement learning tasks.
Societal Impacts and Ethical Issue. Our benchmark has the potential to accelerate the development of deep model fusion algorithms. This can play a crucial role in speeding up the development of machine learning models. By optimizing and combining models more efficiently, we can substantially reduce the carbon footprint associated with training these models, thereby making a positive contribution to environmental sustainability. Given these substantial benefits, the potential negative societal impacts of our work are minimal.
Appendix A Evaluation of Fine-Tuned Single-Task Models
In this section, we describe the experimental setup for fine-tuning the models and present the performance results of the fine-tuned single-task models.
  1. 1.
    CLIP-ViT-B/32 Models: The CLIP models are fine-tuned on eight image classification tasks: SUN397, Cars, RESISC45, EuroSAT, SVHN, GTSRB, MNIST, and DTD. The Adam Optimizer is employed with a fixed learning rate of 1e-5 for a total of 4000 training steps with the batch size of 32. Only the vision encoder is fine-tuned to maintain the model’s open-vocabulary characteristic. The performance of fine-tuned CLIP-ViT-B/32 and CLIP-ViT-L/14 models on the eight image classification tasks is shown in Tables 7 and 8, respectively. In Figure 5, we visualize the cosine similarity matrices of task vectors for CLIP-ViT-B/32 and CLIP-ViT-L/14 models. We note that the task vectors for models from various tasks are nearly orthogonal. This suggests that the knowledge specific to each task resides in distinct directions or subspaces. This finding motivates the exploration of locating subspaces in which the knowledge of different tasks can be merged effectively, as discussed in Tang et al. [50].
  2. 2.
    ResNet-50 Models: We fine-tune ResNet-50 models on three scene understanding tasks: segmentation, depth estimation, and normal estimation using the NYUv2 dataset, each with a learning rate of 1e-4 for 40 epochs, the learning rate is reduced by a factor of 0.5 every 10 epochs. The performance of fine-tuned single-task ResNet-50 models on the NYUv2 dataset is shown in Table 4.
  3. 3.
    GPT-2 Models: GPT2 model fine-tuned on tasks from GLUE benchmark, using a constant learning rate of 5e-5 for 3 epochs. The performance of fine-tuned single-task GPT-2 models on the seven text classification tasks is shown in Table 9.
  4. 4.
    Flan-T5 Models: In this work, we fine-tune Flan-T5-base and Flan-T5-large models on eight text-to-text generation tasks from the GLUE benchmark. The results of LoRA fine-tuned Flan-T5-base and Flan-T5-large models are shown in Tables 10 and 11, respectively.
Table 7: Performance of fine-tuned single-task CLIP-ViT-B/32 models on the eight image classification tasks.
Model SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
Pre-trained 63.2 59.8 60.7 46.0 31.6 32.5 48.2 43.9
SUN397 75.0 47.0 54.3 46.5 28.3 26.4 44.3 41.6
Cars 56.6 78.3 50.9 38.4 30.2 30.6 49.7 41.8
RESISC45 52.0 47.2 95.2 56.9 23.9 24.3 39.7 35.9
EuroSAT 49.0 39.9 33.5 99.0 11.8 22.9 33.8 35.5
SVHN 40.5 36.3 18.9 9.8 97.3 27.3 81.8 23.2
GRSRB 36.9 33.0 20.6 21.3 41.2 98.9 30.9 23.9
MNIST 50.3 40.0 31.3 17.7 50.1 19.3 99.6 30.7
DTD 54.6 51.3 36.8 25.0 28.9 21.8 47.3 79.7
Table 8: Performance of fine-tuned single-task CLIP-ViT-L/14 models on the eight image classification tasks.
Model SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
Pre-trained 68.3 77.8 71.0 58.9 58.4 50.6 76.4 55.5
SUN397 82.8 68.4 58.1 49.9 55.0 46.3 79.5 52.8
Cars 67.8 92.9 68.7 56.4 51.7 47.7 80.5 55.6
RESISC45 65.6 69.0 97.4 64.3 38.3 46.6 77.7 49.9
EuroSAT 65.2 69.0 40.6 99.2 33.4 45.6 73.5 47.1
SVHN 66.5 69.0 54.0 19.7 97.9 48.7 92.2 50.1
GRSRB 63.4 64.8 38.7 19.6 71.0 99.2 75.1 45.8
MNIST 56.1 49.8 53.5 26.6 48.2 33.1 99.8 47.1
DTD 66.8 75.3 65.5 43.7 49.5 45.0 68.5 85.5
(a) CLIP-ViT-B/32 models.
(b) CLIP-ViT-L/14 models.
Figure 5: Cosine similarity matrices of task vectors for CLIP-ViT-B/32 and CLIP-ViT-L/14 models.
Table 9: Performance of fine-tuned single-task GPT-2 models on the seven text classification tasks.
Model CoLA MNLI MRPC QNLI QQP RTE SST-2
CoLA 76.8 32.8 68.4 50.4 39.2 48.0 51.0
MNLI 59.5 82.1 33.8 46.5 24.9 57.4 40.5
MRPC 30.8 25.9 80.4 47.1 65.9 49.1 49.1
QNLI 58.7 38.9 30.6 88.3 39.9 48.7 47.0
QQP 31.4 25.7 62.3 45.0 89.6 49.1 49.1
RTE 52.8 47.7 37.5 53.5 33.7 65.3 54.9
SST-2 51.8 32.9 40.2 49.8 56.8 44.4 91.2
Table 10: Performance of LoRA fine-tuned Flan-T5-Base models on the eight text-to-text generation tasks from the GLUE benchmark.
Model CoLA MNLI MRPC QNLI QQP RTE SST2 STSB
Pre-trained 69.1 56.5 76.2 88.4 82.1 80.1 91.2 62.2
CoLA 69.1 39.9 75.2 89.1 81.1 81.9 90.7 54.0
MNLI 69.4 82.7 73.8 89.3 82.0 79.4 90.9 68.1
MRPC 64.0 44.9 85.5 82.6 81.0 69.0 88.6 73.6
QNLI 68.9 52.7 76.7 90.9 82.8 79.8 91.5 68.9
QQP 65.0 54.6 75.7 89.0 84.0 81.6 90.7 75.3
RTE 64.9 51.8 69.4 89.2 79.8 84.5 90.6 70.1
SST2 68.3 56.6 76.0 88.5 83.4 79.8 92.9 62.6
STSB 65.7 1.7 67.4 89.3 80.1 79.8 90.8 87.4
Table 11: Performance of LoRA fine-tuned Flan-T5-Large models on the eight text-to-text generation tasks from the GLUE benchmark.
Model CoLA MNLI MRPC QNLI QQP RTE SST2 STSB
Pre-trained 73.7 56.6 82.4 91.1 85.5 85.6 94.3 87.5
CoLA 80.2 53.9 81.4 90.8 84.5 84.1 93.9 87.1
MNLI 73.7 88.5 77.9 92.4 85.2 87.7 94.4 86.7
MRPC 75.6 52.6 89.2 92.6 84.4 86.3 94.3 86.3
QNLI 73.5 54.5 82.8 94.4 85.8 85.2 93.7 87.1
QQP 74.0 53.8 82.8 92.5 87.2 85.6 94.5 88.3
RTE 75.6 57.5 69.9 92.8 83.8 91.7 94.6 86.0
SST2 73.6 55.3 82.1 91.6 85.5 85.2 95.2 86.9
STSB 73.4 39.3 82.1 92.6 86.1 83.4 94.0 90.9
Based on the performance metrics detailed in these tables, we observe that the fine-tuned models demonstrate high accuracy on specific tasks. This observation holds true across various model architectures and task domains, indicating the effectiveness of the fine-tuning process in adapting pre-trained models to excel in particular applications.
What’s more, fine-tuning a model on one task can lead to both positive and negative transfer effects on other tasks. Positive transfer occurs when the knowledge gained from fine-tuning on one task enhances the model’s performance on another task, while negative transfer arises when the fine-tuning process on one task hinders the model’s ability to perform well on other tasks.
Appendix B Multi-Task Model Fusion
Figure 6: Merging ResNet-50 models on three scene understanding tasks: segmentation, depth estimation, and normal estimation. Where the backbones are merged and the heads are kept separate.
Table 12: Multi-task performance when merging GPT-2 models on seven text classification tasks.
Method CoLA MNLI MRPC QNLI QQP RTE SST-2 Avg.
Reference Methods
Fine-tuned (STL) 76.8 82.1 80.4 88.3 89.6 65.3 91.2 82.0
Multi-Task Model Fusion Method
Simple Average [57] 55.0 55.1 51.0 57.6 76.7 44.8 52.5 56.1
Fisher Merging [37] 54.8 58.0 39.5 63.3 81.5 49.1 64.7 58.7
RegMean [28] 61.7 70.4 65.4 69.7 78.8 56.0 79.7 68.8
Task Arithmetic [25] 68.7 68.6 69.6 70.5 81.8 47.3 83.6 70.0
Ties-Merging [62] 68.4 71.4 68.4 69.6 82.4 47.7 81.8 70.0
Table 13: Experimental results of merging Flan-T5-base (LoRA fine-tuned) models on all eight tasks.
Method CoLA MNLI MRPC QNLI QQP RTE SST2 STSB Avg.
Reference Methods
Pre-trained 69.1 56.5 76.2 88.4 82.1 80.1 91.2 62.2 75.7
Individual 69.1 82.7 85.5 90.9 84.0 84.4 92.9 87.4 84.6
Multi-Task Model Fusion Methods
Weight Averaging [57] 69.7 59.7 78.9 90.1 83.8 80.5 91.2 72.0 78.2
Task Arithmetic [25] 68.8 55.2 78.7 89.8 83.7 79.1 91.5 72.4 77.4
Ties-Merging [62] 68.3 56.3 79.4 89.8 83.7 79.4 91.6 71.2 77.5
Layer-wise AdaMerging [63] 69.1 60.3 78.4 90.0 83.6 79.1 91.6 74.1 78.3
Table 14: Multi-task performance when merging CLIP-ViT-L/14 models on all eight tasks.
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
Reference Methods
Pre-trained 68.3 77.8 71.0 58.9 58.4 50.6 76.4 55.5 64.6
Fine-tuned (STL) 82.8 92.9 97.4 99.2 97.9 99.2 99.8 85.5 94.3
Multi-Task Model Fusion Methods
Weight Averaging [57] 72.5 81.5 82.2 90.0 81.6 74.0 96.6 61.8 80.0
Fisher Merging [37] 70.6 79.4 84.1 98.1 74.7 85.0 89.5 61.0 80.3
RegMean [28] 75.3 88.4 90.0 97.1 95.9 92.4 98.5 72.6 88.8
Task Arithmetic [25] 72.0 79.0 80.5 86.0 87.5 83.5 98.0 58.8 80.7
Ties-Merging [62] 74.7 83.3 86.4 91.3 89.7 85.2 97.8 63.9 84.0
task-wise AdaMerging [63] 75.8 80.1 77.2 83.6 68.4 93.5 93.1 69.0 80.1
layer-wise AdaMerging [63] 78.1 90.7 90.8 96.5 94.8 97.5 98.6 81.3 91.0
WEMoE (Model Mixing) [51] 81.5 92.3 96.5 98.8 97.6 99.4 99.6 84.5 93.8
Table 15: Experimental results of merging Flan-T5-large (LoRA fine-tuned) models on all eight tasks.
Method CoLA MNLI MRPC QNLI QQP RTE SST2 STSB Avg.
Reference Methods
Pre-trained 73.7 56.6 82.4 91.1 85.5 85.6 94.3 87.5 82.1
Individual 80.2 88.5 89.2 94.4 87.2 91.7 95.2 90.9 89.6
Multi-Task Model Fusion Methods
Weight Averaging [57] 74.6 84.3 84.1 92.8 86.3 87.4 94.8 88.0 86.5
Task Arithmetic [25] 76.9 85.4 85.3 93.9 85.8 88.1 95.2 87.8 87.3
Ties-Merging [62] 77.1 85.1 86.3 93.9 86.0 87.7 95.1 88.0 87.4
Layer-wise AdaMerging [63] 76.7 87.6 84.8 93.8 85.9 88.1 95.2 88.6 87.6
We begin by comparing the performance of multi-task model fusion algorithms on various tasks using different models. These experiments provide insights into the effectiveness of different fusion methods in improving the performance of multi-task models. We evaluate the performance of multi-task model fusion algorithms on image classification tasks using CLIP models, scene understanding tasks using ResNet-50 models, text classification tasks using GPT-2 models, and text-to-text generation tasks using Flan-T5 models.
  1. 1.
    Image Classification Tasks with CLIP Models: We utilize the CLIP-ViT-B/32 and CLIP-ViT-L/14 models from the HuggingFace Library [24]. The results of merging CLIP-ViT-B/32 models on all eight tasks are provided in Table 3 and Figure 4. The results of CLIP-ViT-L/14 models are shown in Table 14.
  2. 2.
    Scene Understanding Tasks with ResNet-50 Models: We use the NYUv2 dataset and ResNet-50 models for segmentation, depth estimation, and normal estimation tasks. In Figure 6, we illustrate the process of merging ResNet-50 models on these tasks, where the backbones are merged, and the heads are copied separately. The results of merging ResNet-50 models on these tasks are shown in Table 4.
  3. 3.
    Text Classification Tasks with GPT-2 Models: The results of merging GPT-2 models on seven text classification tasks are shown in Table 12.
  4. 4.
    Text-to-Text Generation Tasks with Flan-T5 Models: For LoRA fine-tuned Flan-T5-base and Flan-T5-large models, we merge and unload the LoRA adapters before performing multi-task model fusion. The results of merging Flan-T5-base and Flan-T5-large models on all eight tasks are shown in Tables 13 and 15, respectively.
In the above mentioned tables, we compare the performance of different multi-task model fusion algorithms on various tasks. The results of pre-trained models, fine-tuned models, and traditional multi-task learning (MTL) are provided as reference methods.
Appendix C Generalization Experiments
Table 16: Generalization results on two unseen tasks when merging ViT-B/32 models on six tasks.
Method Seen Tasks Unseen Tasks
SUN397 Cars GTSRB EuroSAT DTD MNIST Avg. RESISC45 SVHN Avg.
Pre-trained 63.2 59.9 30.4 45.6 43.9 47.6 48.4 60.6 23.5 40.1
Fisher Merging [37] 68.1 67.4 67.2 86.4 58.6 81.6 71.5 60.2 42.5 51.3
RegMean [28] 69.4 70.5 86.9 97.0 67.1 98.3 81.5 50.2 51.5 50.8
Task Arithmetic [25] 65.2 63.6 76.1 87.1 56.4 94.2 73.8 52.4 45.2 48.8
Ties-Merging [62] 68.2 65.9 70.0 81.2 56.0 89.0 71.7 60.3 47.3 53.8
AdaMerging [63] 69.8 72.4 95.5 95.1 70.7 98.1 83.6 48.7 60.7 54.7
WEMoE [51] 74.3 78.1 98.8 98.7 75.1 99.5 87.4 47.3 51.3 49.3
For the generalization experiments, we assess the performance of multi-task model fusion algorithms on two unseen tasks after merging ViT-B/32 models trained on six tasks. The performance of various multi-task model fusion methods, including Fisher Merging [37], RegMean [28], Task Arithmetic [25], Ties-Merging [62], AdaMerging [63], and WEMoE [51], is compared across both the seen tasks and unseen tasks.
Specifically, we conduct two sets of generalization experiments using the CLIP-ViT-B/32 models:
  • In the first set, we merge models trained on six tasks (SUN397, Cars, RESISC45, DTD, SVHN, GTSRB) and evaluate the fused model on the unseen tasks (MNIST, EuroSAT). The results are shown in Table 5.
  • In the second set of experiments, we merge models trained six tasks (SUN397, Cars, GTSRB, EuroSAT, DTD, MNIST) and evaluate the fused model on the unseen tasks (RESISC45, SVHN). The results are shown in Table 16.
By conducting these two sets of generalization experiments, we aim to gain a comprehensive understanding of how the CLIP-ViT-B/32 models, when fused with knowledge from different task combinations, can perform on various unseen tasks. From these experimental results, we can observe instances of negative transfer when evaluating the fused CLIP models on unseen tasks. Here, negative transfer occurs when the knowledge gained from fine-tuning on a set of tasks hinders the model’s performance on new, unseen tasks. In other words, the model’s ability to generalize and adapt to novel challenges is compromised due to the specific knowledge acquired during the fine-tuning process. The presence of negative transfer in these experiments highlights the challenges and limitations of model fusion and generalization in the context of the CLIP-ViT-B/32 models. Several factors can contribute to negative transfer, such as:
  • Task dissimilarity: If the unseen tasks are significantly different from the tasks used for fine-tuning, the learned representations may not be directly applicable, leading to performance degradation.
  • Overspecialization: Fine-tuning on a specific set of tasks may cause the model to overfit to task-specific features and patterns, reducing its ability to generalize to new tasks.
  • Interference between tasks: When merging knowledge from multiple tasks, there may be conflicts or interference between the learned representations, hindering the model’s ability to adapt to unseen tasks effectively.
To mitigate the negative transfer and improve the generalization ability of merged models, several strategies can be explored, such as:
  • Task selection: Carefully selecting tasks that are more similar or complementary to the target unseen tasks can help reduce the risk of negative transfer. This is adapted in Wu et al. [58], where the Fisher information matrix is computed for a proxy metric for task similarity.
  • Regularization techniques: Applying regularization methods, such as weight decay or dropout, during the fine-tuning process may help prevent overfitting and promote better generalization.
Appendix D Prompt-Based Text-to-Text Generation
This section details the prompt templates employed for each of the eight text-to-text generation tasks from the GLUE benchmark., see Section 3.3 for more details. Within each task, we provide the format of the input text, and the corresponding target text mapping. These templates are crucial in fine-tuning the Flan-T5 models for generating appropriate text outputs tailored to each specific task.
  • CoLA:
    • Input Text: "Indicate if the following sentence is grammatically correct or not: "sentence". Answer ‘acceptable’ or ‘unacceptable’."
    • Target Text:
        0: "unacceptable"
        1: "acceptable"
  • MNLI:
    • Input Text: "Does the premise: ‘premise’ logically imply, contradict, or is neutral to the hypothesis: ‘hypothesis’? Answer with ‘entailment’, ‘contradiction’, or ‘neutral’."
    • Target Text:
        0: "entailment"
        1: "neutral"
        2: "contradiction"
  • MRPC:
    • Input Text: "Are the following sentences ‘sentence1’ and ‘sentence2’ conveying the same meaning? Answer with ‘yes’ or ‘no’."
    • Target Text:
        0: "no"
        1: "yes"
  • QNLI:
    • Input Text: "Given the context: ‘sentence’, does the question ‘question’ have an answer based on the information provided? Answer with ‘yes’ or ‘no’."
    • Target Text:
        0: "yes"
        1: "no"
  • QQP:
    • Input Text: "Do the questions ‘question1’ and ‘question2’ have the same intent? Answer with ‘yes’ or ‘no’."
    • Target Text:
        0: "no"
        1: "yes"
  • RTE:
    • Input Text: "Does the text: ‘sentence1’ entail that ‘sentence2’ is true? Provide ‘yes’ or ‘no’."
    • Target Text:
        0: "yes"
        1: "no"
  • SST-2:
    • Input Text: "Given the sentence ‘sentence’, determine the sentiment. Is it positive or negative?"
    • Target Text:
        0: "negative"
        1: "positive"
  • STSB:
    • Input Text: "Consider the sentences ‘sentence1’ and ‘sentence2’. On a scale from 1 (completely different) to 5 (completely similar), rate the similarity."
    • Target Text: ":.1f", parse to float with one decimal place

### v4

@@
-FusionBench: A Comprehensive Benchmark of Deep Model Fusion
+FusionBench: A Unified Library and Comprehensive Benchmark for Deep Model Fusion
 Abstract
-Deep model fusion is an emerging technique that unifies the predictions or parameters of several deep neural networks into a single model in a cost-effective and data-efficient manner.
-This enables the unified model to take advantage of the original models’ strengths, potentially exceeding their performance.
-Although a variety of deep model fusion techniques have been introduced, their evaluations tend to be inconsistent and often inadequate to validate their effectiveness and robustness against distribution shifts.
-To address this issue, we introduce FusionBench, which is the first comprehensive benchmark dedicated to deep model fusion.
-FusionBench covers a wide range of tasks, including open-vocabulary image classification, text classification, and text-to-text generation.
-Each category includes up to eight tasks with corresponding task-specific models, featuring both full fine-tuning and LoRA fine-tuning, as well as models of different sizes, to ensure fair and balanced comparisons of various multi-task model fusion techniques across different tasks, model scales, and fine-tuning strategies.
-We implement and evaluate a broad spectrum of deep model fusion techniques.
-These techniques range from model ensemble methods, which combine the predictions to improve the overall performance, to model merging, which integrates different models into a single one, and model mixing methods, which upscale or recombine the components of the original models.
-FusionBench now contains 26 distinct tasks, 74 fine-tuned models, and 16 fusion techniques, and we are committed to consistently expanding the benchmark with more tasks, models, and fusion techniques.
-In addition, we offer a well-documented set of resources and guidelines to aid researchers in understanding and replicating the benchmark results.
-This includes detailed documentation, code examples, and tutorials, making FusionBench a user-friendly and accessible platform for both beginners and experienced researchers.
-The FusionBench homepage is available at https://github.com/tanganke/fusion_bench.
+Deep model fusion is an emerging technique that unifies the predictions or parameters of several deep neural networks into a single better-performing model in a cost-effective and data-efficient manner.
+Although a variety of deep model fusion techniques have been introduced, their evaluations tend to be inconsistent and often inadequate to validate their effectiveness and robustness.
+We present FusionBench, the first benchmark and a unified library designed specifically for deep model fusion.
+Our benchmark consists of multiple tasks, each with different settings of models and datasets.
+This variety allows us to compare fusion methods across different scenarios and model scales.
+Additionally, FusionBench serves as a unified library for easy implementation and testing of new fusion techniques.
+FusionBench is open source and actively maintained, with community contributions encouraged.
+[Uncaptioned image] Code Repository: https://github.com/tanganke/fusion_bench
+[Uncaptioned image] Online Documentation: https://tanganke.github.io/fusion_bench
+[Uncaptioned image] HuggingFace Collections (models & datasets): https://huggingface.co/tanganke
+Keywords: Deep Model Fusion, Model Merging, Multi-Task Learning, Benchmark
 1 Introduction
-In recent years, a new paradigm called “learn from model” has emerged in the field of deep learning, which focuses on leveraging the knowledge embedded in existing models to develop new ones [67].
-This paradigm has been widely adopted in various scenarios, such as model tuning [18, 8], model distillation [21], model pruning [16, 3], model editing [38, 66], and so on.
+In recent years, a new paradigm called “learn from model” has emerged in the field of deep learning, which focuses on leveraging the knowledge embedded in existing models to develop new ones (zheng2023learn).
+This paradigm has been widely adopted in various scenarios, such as model tuning (he2022masked; chung2024scaling), model distillation (hinton2015distilling), model pruning (han2015deep; asif2020ensemble), model editing (mitchell2021fast; zhang2024comprehensive), and so on.
 Among these methods, deep model fusion is particularly appealing.
 It merges the parameters or predictions of multiple models to create a more robust and efficient unified model.
-Due to its effectiveness and scalability, many new techniques for deep model fusion have recently been proposed [34].
+Due to its effectiveness and scalability, many new techniques for deep model fusion have recently been proposed (li2023deep).
 Deep model fusion offers both scalability and data efficiency by utilizing the knowledge embedded in pre-existing models, rather than requiring training from scratch.
@@
 Despite its potential, the evaluation of deep model fusion techniques often suffers from inconsistency and inadequacy.
-Standardized assessments are lacking, making it challenging to verify their effectiveness and robustness.
+Current libraries like MergeKit (goddard2024arcee) are primarily focused only on large language model merging and do not provide evaluation frameworks.
+Standardized assessments across different tasks and models are lacking, making it challenging to evaluate the effectiveness and robustness of deep model fusion techniques.
 The potential reasons for this inconsistency include the rapid development of new techniques, the absence of standardized tasks and models, and the variety of settings (such as different fine-tuning strategies).
 Additionally, challenges in implementing or replicating prior work contribute to these inconsistencies.
+A more detailed comparison with existing model fusion methods and toolkits is provided in Section A.
 To tackle these challenges, we develop the first comprehensive benchmark dedicated to deep model fusion, called FusionBench.
-FusionBench is built to be a modular and extensible platform, comprising three core modules: the Algorithm Module, the Model Pool Module, and the Task Pool Module.
-Each module is configurable, allowing users to easily customize and manage the benchmark settings.
-In summary, our main contributions are four-fold:
-  1. 1.
-    An extensive codebase and configurable interface: Our codebase is structured around three core modules: the algorithm module, the model pool module, and the task pool module.
-Moreover, to facilitate ease of use and flexibility, we provide a unified command line interface with YAML configuration options for effortless customization.
-  2. 2.
-    Comprehensive evaluations: We conduct thorough evaluations of the deep model fusion techniques, assessing their performance across a variety of benchmarks and settings.
-  3. 3.
-    Extensive analysis and insights: We provide detailed analysis and insights based on the evaluation results, revealing key factors that influence their effectiveness.
-This includes identifying best practices, exploring the impact of fine-tuning strategies, and offering recommendations for future research.
-  4. 4.
-    User-friendly resources and guidelines: We offer a well-documented set of resources and guidelines to aid researchers in understanding and replicating the benchmark results.
-This includes detailed documentation, code examples, and tutorials, making FusionBench a user-friendly and accessible platform for both beginners and experienced researchers.
-Documentation is available at https://tanganke.github.io/fusion_bench/.
-2 Related Work
+FusionBench is designed as a modular and flexible framework with three primary components: the Algorithm Module, the Model Pool Module, and the Task Pool Module as shown in Figure 1.
+Each component offers full configurability, enabling straightforward integration of novel algorithms, models, and evaluation tasks.
+Through Hydra-based configuration management, all experimental setups are streamlined via a single unified interface.
+Furthermore, FusionBench provides extensive documentation, hands-on tutorials, illustrative code samples, and ready-to-use model collections.
+Figure 1: The general framework of the modularized FusionBench codebase.
+2 FusionBench
+FusionBench is available for installation through two simple methods: using pip with ‘pip install fusion-bench’ or cloning the repository directly with ‘git clone’.
+The general framework of the modularized FusionBench codebase is shown in Figure 1, which consists of three primary elements: Algorithm Module, Model Pool Module, and Task Pool Module.
+In Section 2.1, we introduce the three major components, which is designed to be flexible and modular, allowing users to easily run experiments and evaluate the performance of model fusion algorithms.
+In Section 2.2, we introduce the implemented model fusion algorithms.
+Finally, in Section 2.3, we discuss the documentation and tutorials provided to help users understand the benchmark and effectively use the codebase.
+In Appendix C, we provide a flowchart to illustrate the process of running experiments and evaluating the merged models.
+Main Program Pipeline.
+The typical pipeline consists of the following steps: (1) Users configure the experiment by selecting a fusion algorithm, models to merge, and optional evaluation tasks.
+(2) The system loads the specified models, algorithm, and evaluation tasks.
+(3) The algorithm merges the selected models, with an option to save the resulting model.
+(4) When evaluation tasks are specified, the merged model is tested and results are reported.
+2.1 Components
+As shown in Figure 1, the codebase is composed of three primary components: Algorithm Module, Model Pool Module, and Task Pool Module, which are responsible for implementing the model fusion algorithms, managing the models and datasets used during the algorithm execution, and managing the tasks and evaluation metrics to be evaluated, respectively.
+Additionally, we provide a unified command line interface (CLI) ‘fusion_bench’ to facilitate the use of the codebase and to enable users to easily run experiments.
+  • Algorithm Module is the core component, which contains the implementation of all model fusion algorithms.
+Each algorithm is implemented as a separate Python class inheriting from the base class fusion_bench.BaseAlgorithm.
+The algorithm classes are flexible - they can be used through our CLI or in custom code.
+  • Model Pool Module provides a unified interface for managing models and datasets.
+  • Task Pool Module evaluates model performance on different tasks, where each task combines a dataset with specific evaluation metrics.
+2.2 Implemented Algorithms
+Table 1: Implemented algorithms in FusionBench.
+Method Requirement
+Model Ensemble Methods
+Simple Ensemble (sagi2018ensemble) -
+Weighted Ensemble (sagi2018ensemble) hyperparameter search
+Max-Model Predictor (wu2019heterogeneous) -
+Model Merging Methods
+Simple Average / Modelsoups (wortsman2022model) -
+Weighted Average (matena2022merging) hyperparameter search
+Fisher Merging (matena2022merging) compute weights on labeled data
+RegMean (jin2022dataless) compute weights on labeled data
+RegMean++ (nguyen2025regmean++) compute weights on labeled data
+Concrete Subspace (tang2023concrete) test-time adaptation training
+Task Arithmetic (ilharco2022editing) hyperparameter search
+Ties-Merging (yadav2023resolving) hyperparameter search
+Task-Wise AdaMerging (yang2023adamerging) test-time adaptation training
+Layer-Wise AdaMerging (yang2023adamerging) test-time adaptation training
+Representation Surgery (yang2024representation) unlabeled data samples
+TALL mask (wang2024localizing) -
+Task Singular Vectors (gargiulo2024task) -
+Isotropic Merging (marczak2025no) -
+OPCM (tang2025merging) -
+FW-Merging (chen2025fw) labeled data samples
+RanDeS (zhou2025randes) -
+Model Mixing Methods
+MoE-based Upscaling (komatsuzaki2022sparse) pre-training to recover performance
+MoE-based Merging (komatsuzaki2022sparse) training on the combined model
+Depth Upscaling (kim2023solar) pre-training to recover performance
+Model Recombination (hu2023fedmr) training on the combined model
+Weight-Ensemble MoE (tang2024merging) test-time adaptation training, vision tasks
+WE-MoE V2 (shen2024efficient) test-time adaptation training, vision tasks
+Pareto-Driven Merging (tang2024towards) training datasets
+SMILE Upscaling (tang2024smile) -
+Model Compression Methods
+BitDelta (liu2024bitdelta) calibration data
+Magnitude Pruning (frankle2020linear) -
+Wanda Pruning (sun2023simple) calibration data
+SparseGPT Pruning (frantar2023sparsegpt) calibration data
+Expert Pruning (lu2024not) calibration data
+Our benchmark incorporates a comprehensive collection of cutting-edge model fusion algorithms, spanning three key categories: ensemble methods, merging methods, and mixing methods.
+Additionally, we have recently extended FusionBench to include model compression methods, which, while not strictly fusion techniques, are complementary approaches that can be applied to reduce the computational overhead of models or compress individual models before and after fusion.
+Additionally, we evaluated their prominence in the research literature and their practical utility.
+The complete listing of these implemented algorithms can be found in Table 1 and FusionBench is under active development and will be updated to include the latest and most effective model fusion algorithms.
+2.3 Online Documentation and Tutorials
+We provide detailed documentation and tutorials on our project website to help users get started.
+These resources cover the basics of model fusion, how to run experiments, and how to evaluate results.
+We also include example pre-run results showing how different fusion algorithms perform across various tasks.
+3 Conclusion and Discussion
+FusionBench offers a robust framework for evaluating deep model fusion algorithms and a unified library for developing new ones.
+Its scalable and extensible architecture simplifies the creation of deep model fusion techniques.
+Furthermore, FusionBench provides curated datasets and models to ensure fair comparisons.
+With comprehensive documentation and tutorials, it is accessible to both beginners and experienced researchers.
+We encourage the community to use this benchmark to advance the field of deep model fusion.
+Acknowledgments and Disclosure of Funding
+This work is supported by the National Natural Science Foundation of China (Grant No.
+62225113, U23A20318, U2336211, 62576364 and 62276195), the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (Grant No.
+JYB2025XDXM704), the Shenzhen Basic Research Project (Natural Science Foundation) Basic Research Key Project (NO.
+JCYJ20241202124430041), the Foundation for Innovative Research Groups of Hubei Province (Grant No.
+2024AFA017) and the Science and Technology Major Project of Hubei Province (Grant No.
+2024BAB046).
+Dr.
+Tao’s research is partially supported by NTU RSR and Start Up Grants.
+The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
+The appendix is organized into several sections, each providing additional insights and details related to different aspects of the main work.
+Appendix A Related Work of Deep Model Fusion
 Figure 2: A taxonomy of deep model fusion techniques.
-Since deep model fusion is a relatively new research area, there is currently no standardized taxonomy.
+Deep model fusion, often referred to in the literature simply as “model fusion”, aims to combine the outputs or parameters of multiple deep neural networks into a unified model with enhanced performance, efficiency, or robustness.
 Different researchers may categorize these techniques in various ways based on their understanding and points of view.
@@
 A visualization of the taxonomy is shown in Figure 2.
-Model Ensemble methods combine the predictions of multiple models to improve the overall performance of a machine learning system [45], where the collective knowledge of multiple models is often more accurate and reliable than that of any individual model.
-Mathematically, given a set of $N$ models $\{f_{1},f_{2},\dots,f_{N}\}$, which can be homogeneous or heterogeneous, we use their predictions to obtain a global prediction $y=\mathcal{A}(x;f_{1},f_{2},\dots,f_{N};w)$, where $\mathcal{A}$ is an ensemble algorithm and $w$ are the algorithmic parameters.
-Each model $f_{i}$ can also be associated with a specification to indicate its weight or importance in the ensemble [41, 68, 60, 49].
+A.1 A Taxonomy of Deep Model Fusion Techniques
+Model Ensemble methods combine the predictions of multiple models to improve the overall performance of a machine learning system (sagi2018ensemble), where the collective knowledge is often more accurate and reliable than that of any individual model.
+Mathematically, given a set of $N$ models $\{f_{1},f_{2},\dots,f_{N}\}$, which can be homogeneous or heterogeneous, we use their predictions to obtain a global prediction $y=\mathcal{A}_{ensemble}(x;f_{1},f_{2},\dots,f_{N};w)$, where $\mathcal{A}_{ensemble}$ is an ensemble algorithm and $w$ are the algorithmic parameters.
+Each model $f_{i}$ can also be associated with a specification to indicate its weight or importance in the ensemble (pathak2010multiparty; zhouLearnwareFutureMachine2016; wu2021model; tangImprovingHeterogeneousModel2023a).
 Ensemble methods are widely used and effective in improving performance but are often expensive to use and manage.
-Recently, there have also been studies exploring efficient methods for model ensembles [55, 5, 2].
-Model Merging methods integrate the parameters of multiple models into a single model, which can be more efficient and scalable model fusion techniques.
-Mathematically, given a set of $N$ isomorphic models $\{f_{i}(\cdot;\theta_{i})\}_{i=1}^{N}$, each parameterized with $\theta_{i}$, we merge them into a single model with parameters $\theta=\mathcal{A}(\theta_{1},\theta_{2},\dots,\theta_{N};w)$, where $\mathcal{A}$ is a merging algorithm and $w$ are the algorithmic parameters.
+Recent research has also investigated efficient techniques for model ensembles (wen2020batchensemble; chen2023split; allingham2021sparse).
+Model Merging methods integrate the parameters of multiple models into a unified model, enhancing efficiency in terms of inference cost and storage, and enabling scalable model fusion.
+Given a set of $N$ isomorphic models $\{f_{i}(\cdot;\theta_{i})\}_{i=1}^{N}$, each parameterized with $\theta_{i}$, we merge them into a single model with parameters $\theta=\mathcal{A}_{merging}(\theta_{1},\theta_{2},\dots,\theta_{N};w)$, where $\mathcal{A}_{merging}:\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{d}$ is a merging algorithm and $w$ are the algorithmic parameters.
 The merged model can be expressed as $f(\cdot;\theta)$.
-This method can be implemented through linear interpolation in parameter space [57, 25, 62, 37, 65, 7, 44, 15, 36], leveraging mode connectivity [10, 11, 4, 12, 42], aligning features or parameters [35, 1, 28, 48, 13, 26], and ensemble distillation [52, 53].
-Model merging methods are often performed in a data-efficient manner, the algorithmic parameters $w$ can also be learned during test time via test-time adaptation training or meta-learning for a more seamless merging [63, 50].
-Model Mixing methods fuse the components of multiple models to create a new model, which can be more flexible and adaptive than the original models.
-Mathematically, given a set of $N$ models $\{f_{i}(\cdot;\theta_{i})\}_{i=1}^{N}$, each parameterized with $\theta_{i}$, we mix their components to obtain a new model with parameters $\Theta=\mathcal{A}(\theta_{1},\theta_{2},\dots,\theta_{N};w)$, where $\mathcal{A}$ is a mixing algorithm and $w$ is the algorithmic parameters.
+This method can be implemented through linear interpolation in parameter space (wortsman2022model; ilharco2022editing; yadav2023resolving; matena2022merging; yu2024language; chronopoulou2023adaptersoup; rame2024rewarded; ortiz2024task; liu2023tangent), leveraging mode connectivity (draxler2018essentially; frankle2020linear; benton2021loss; garipov2018loss; qu2024rethink), aligning features, parameters or gradients (liu2022deep; ainsworthGitReBasinMerging2023; jin2022dataless; tam2024merging; stoica2023zipit; jang2023personalized; daheim2023model; yang2024representation), subspace-based methods (tang2023concrete; wang2024localizing; yi2024safety; zhu2024model; xu2024weight), and ensemble distillation (wan2024knowledge; wan2024fusechat).
+Model merging methods are often performed in a data-efficient manner, the algorithmic parameters $w$ can also be learned during test time via test-time adaptation (TTA) training or meta-learning for a more seamless merging (yang2023adamerging; tang2023concrete).
+Model Mixing methods fuse the components of multiple models to create a new heterogeneous model, which can be more flexible and adaptive than the original models.
+Mathematically, given a set of $N$ models $\{f_{i}(\cdot;\theta_{i})\}_{i=1}^{N}$, each parameterized with $\theta_{i}\in\mathbb{R}^{d}$, we mix their components to obtain a new model with parameters $\Theta=\mathcal{A}_{mixing}(\theta_{1},\theta_{2},\dots,\theta_{N};w)\in\mathbb{R}^{d^{\prime}}$, where $\mathcal{A}_{mixing}:\mathbb{R}^{N\times d}\mapsto\mathbb{R}^{d^{\prime}}$ is a mixing algorithm and $w$ is the algorithmic parameters.
 The mixed model can be expressed as $F(\cdot;\Theta)$, which often has more parameters than the original models, and thus can be more expressive and powerful to capture the underlying patterns in the data.
-Model mixing methods can be implemented through layer recombinations [23, 27], model stitching [33, 39], or upscale to create a Mixture of Experts (MoE)-based sparse model [30, 64, 51].
-Although several model fusion methods have been proposed, benchmarks and unified toolkits are still lacking in this field.
-A recent notable work, MergeKit [14], provides a collection of model fusion techniques specifically designed for merging large language models (LLMs), with a focus on model merging methods and Transformer-based LLMs.
-However, MergeKit’s scope is limited to a specific domain and model architecture.
-In contrast, FusionBench aims to be a comprehensive benchmark that encompasses a wide range of tasks and model types.
-It includes a diverse set of fine-tuned models and tasks to evaluate, making it a more generalized and versatile platform for assessing the performance of different model fusion approaches across various domains and architectures.
-3 Our Benchmark
-Figure 2: The general framework of the modularized FusionBench codebase.
-The general framework of the modularized FusionBench codebase is shown in Figure 2, which consists of three primary elements: Algorithm Module, Model Pool Module, and Task Pool Module.
-In Section 3.1, we introduce the codebase, which is designed to be flexible and modular, allowing users to easily run experiments and evaluate the performance of model fusion algorithms.
-In Section 3.2 and Section 3.3, we introduce the implemented model fusion algorithms and the tasks and models included in FusionBench.
-Finally, in Section 3.4, we discuss the documentation and tutorials provided to help users understand the benchmark and effectively use the codebase.
-3.1 Codebase
-We’ve constructed a flexible and modular codebase, which serves as the foundation for FusionBench.
-As shown in Figure 2, the codebase is composed of three primary elements: Algorithm Module, Model Pool Module, and Task Pool Module, which are responsible for implementing the model fusion algorithms, managing the models to be fused, and managing the tasks to be evaluated, respectively.
-Additionally, we provide a command line interface (CLI) to facilitate the use of the codebase and to enable users to easily run experiments and evaluate the performance of model fusion algorithms.
-Algorithm Module is the core component of the codebase, which contains the implementation of various model fusion algorithms.
-Each algorithm is implemented as a separate Python class, which inherits from the base class ModelFusionAlgorithm.
-The algorithm classes are designed to be configurable and independently callable, allowing users to easily instantiate and set up the algorithms through our CLI or by directly invoking the Python classes in their own code.
-Model Pool Module is responsible for managing the models to be fused.
-It offers a unified interface for loading the pre-trained model and fine-tuned models.
-The module is designed to be extensible, allowing users to easily add support for new model architectures and add their own models to the pool.
-Each model in the pool can also be associated with metadata to meet the requirements of specific model fusion algorithms, such as the test dataset for test-time adaptation training.
-Task Pool Module is responsible for managing the tasks to be evaluated.
-Each task comprises a dataset and a set of evaluation metrics, which are defined in the YAML configuration file.
-This module offers a unified interface for loading tasks and assessing the performance of model fusion algorithms on these tasks.
-Users can effortlessly add support for new task types and evaluation metrics, or add new tasks of the same type but with different datasets.
-3.2 Implemented Algorithms
-Table 1: Implemented model fusion algorithms in FusionBench.
-Category Method Requirement
-Model Ensemble Simple Ensemble [45] -
-Weighted Ensemble [45] hyperparameter search
-Max-Model Predictor [59] -
-Model Merging Simple Average / Modelsoups [57] -
-Weighted Average [37] hyperparameter search
-Fisher Merging [37] compute weights on labeled data
-RegMean [28] compute weights on labeled data
-Task Arithmetic [25] hyperparameter search
-Ties-Merging [62] hyperparameter search
-Task-Wise AdaMerging [63] test-time adaptation training
-Layer-Wise AdaMerging [63] test-time adaptation training
-Model Mixing Depth Upscaling [29] pre-training to recover performance
-MoE-based Upscaling [30] pre-training to recover performance
-MoE-based Merging [30] training on the combined model
-Weight-Ensemble MoE [51] test-time adaptation training, vision tasks
-Model Recombination [23] training on the combined model
-In our benchmark, we have implemented 16 model fusion algorithms as the initial set.
-This includes 3 model ensemble methods, 8 model merging methods, and 5 model mixing methods.
-Our primary selection criterion for choosing among various algorithms is their applicability and effectiveness within the realm of deep learning architectures We have also considered the popularity of the algorithms in the literature and their practical applicability, such as their potential use in large-scale language models.
-We list the implemented algorithms in Table 1.
-As shown in Table 1, we implemented three kinds of model fusion algorithms.
-A brief introduction and formal definition of our taxonomy are provided in Section 2.
-Model ensemble methods are effective at enhancing the performance of a machine learning system, but they are computationally expensive to infer.
-Model merging methods aim to integrate the advantages of individual models, making them popular in multi-task model fusion and auxiliary learning.
-In these scenarios, multiple single-task models are merged to construct a multi-task model, or models focused on auxiliary tasks are combined to boost the performance of a primary task.
-Model mixing methods are frequently used to scale up a pre-trained model to a larger size or to combine multiple models into a new one.
-Consequently, model mixing methods often necessitate additional training after the fusion process.
-3.3 Tasks and Models
-Table 2: Tasks and models included in FusionBench for evaluating multi-task model fusion algorithms.
-Domain Task Type Datasets Models
-Computer Vision Image classification (8 tasks) SUN397 [61], Stanford Cars [31], RESISC45 [6], EuroSAT [19], SVHN [40], GTSRB [47], MNIST [32], DTD [9] 8$\times$×CLIP-ViT-B/32 (Transformers [56]), 8$\times$×CLIP-ViT-B/32 (OpenCLIP [24]), 8$\times$×CLIP-ViT-B/16 (OpenCLIP [24]), 8$\times$×CLIP-ViT-L/14 (Transformers [56]), 8$\times$×CLIP-ViT-L/14 (OpenCLIP [24])
-Sence Understanding (3 tasks) NYUv2 [46] 3$\times$×Resnet-50 models
-Natural Language Processing Text classification (7 tasks) CoLA, MNLI, MRPC, QNLI, QQP, RTE, and SST-2 [54] 7$\times$×GPT-2
-Text-to-text generation (8 tasks) CoLA, MNLI, MRPC, QNLI, QQP, RTE, SST-2, and STSB [54] 16$\times$×Flan-T5-Base (w/ & w/o LoRA), 8$\times$×Flan-T5-Large (w/ LoRA)
-(a) Clean
-(b) Motion
-(c) Impulse
-(d) Gaussian
-(e) Pixelate
-(f) Spatter
-(g) Contrast
-(h) JPEG
-Figure 3: Here are eight instances of distorted images from the Stanford Cars dataset. We use these distortions to assess the robustness and generalization capacity of the merged model as in [63, 51].
-Model fusion is a versatile technique that can be applied across various machine learning tasks at different stages of model development.
-In FusionBench, we specifically provide a diverse array of tasks and corresponding fine-tuned models to ensure a fair and comprehensive evaluation of multi-task model fusion algorithms.
-We have selected tasks from the domains of computer vision and natural language processing, as these are the most popular and extensively studied areas in deep learning research.
-The tasks included in our benchmark are open-vocabulary image classification, text classification, and text-to-text generation.
-We list these tasks and models in Table 2.
-We make them publicly available to facilitate reproducibility and further research at HuggingFace  [^1] Access to datasets and fine-tuned models at https://huggingface.co/tanganke.
-Open-vocabulary image classification is a challenging task that requires models to classify images into a large number of categories.
-We have selected eight image classification datasets, including SUN397 [61], Stanford Cars [31], RESISC45 [6], EuroSAT [19], SVHN [40], GTSRB [47], MNIST [32], DTD [9].
-These datasets cover a wide range of image classification tasks, including object recognition, satellite image classification, and texture classification.
-We fine-tuned two CLIP-ViT models, CLIP-ViT-B/32 and CLIP-ViT-L/14, on these datasets.
-We report accuracy as the evaluation metric for these tasks.
-Specifically, to assess the robustness of multi-task model fusion algorithms, particularly those needing test-time adaptation training, we adopt the techniques recommended by Hendrycks and Dietterich [20] to create corrupted versions of the test set for Cars, EuroSAT, RESISC45, and GTSRB.
-These corruptions are designed to simulate common image corruptions in real-world scenarios, including motion blur, impulse noise, Gaussian noise, pixelation, spatter, contrast adjustments, and JPEG compression.
-Scene understanding tasks are performed using the NYUv2 [46] dataset, which consists of RGB-D images and includes three tasks: 13-class segmentation, depth estimation, and surface normal estimation.
-We fine-tuned ResNet-50 models [17] as the backbone for our experiments.
-The initial weights for these models were pre-trained on the ImageNet dataset.
-We then adapted them to the specific tasks of segmentation, depth estimation, and surface normal estimation.
-Text classification is a fundamental task in natural language processing that involves categorizing text data into predefined classes.
-We have selected seven text classification tasks from the General Language Understanding Evaluation (GLUE) benchmark [54], including CoLA, MNLI, MRPC, QNLI, QQP, RTE, and SST-2.
-We fine-tuned GPT-2 models on these seven tasks, each with a different head for classification [43].
-We report accuracy as the evaluation metric.
-Text-to-text generation is more challenging than text classification, as it requires models to generate appropriate text outputs instead of projecting the hidden representations to logits.
-Similar to text classification, we have selected eight text-to-text generation tasks from the GLUE benchmark, including CoLA, MNLI, MRPC, QNLI, QQP, RTE, SST-2, and STSB.
-We fine-tuned Flan-T5 models on these tasks, with and without the LoRA adaptation [22].
-The prompt templates for these tasks are provided in Appendix D.
-We report Spearman’s $\rho$ for STSB and exact match accuracy for other tasks.
-In particular for STSB, if text outputs can’t be parsed as numerical values, we report the $\rho$ as 0.
-3.4 Documentation and Tutorials
-Documentation and tutorials are essential for beginners to understand the methodology behind the benchmark, to reproduce the experiments, and to effectively use the codebase.
-To this end, we offer comprehensive documentation and tutorials on the project homepage, which guide users through the fundamentals of model fusion, the steps to run experiments, and the procedures for evaluating the performance of model fusion algorithms.
-Additionally, we present some experimental results to shed light on the performance of different model fusion algorithms across various tasks.
-4 Evaluation and Analysis
-In this section, we evaluate the performance of multi-task model fusion algorithms on a variety of tasks, as well as analyze the generalization and robustness of these algorithms.
-We also provide an ablation study to investigate the impact of hyperparameter selection.
-Most of the experiments are conducted with a single NVIDIA RTX 3090 GPU with 24GB memory.
-4.1 Experimental Setup
-In this section, we conduct a series of multi-task model fusion experiments on image classification tasks, text classification tasks, and text-to-text generation tasks to evaluate the performance of multi-task model fusion algorithms.
-These tasks are chosen to cover a wide range of NLP and CV tasks, as described in Section 3.3.
-Table 2 provides a summary of the tasks and models used in our experiments.
-4.2 Multi-Task Model Fusion
-Figure 4: Illustration of multi-task model fusion.
-Table 3: Multi-task performance when merging CLIP-ViT-B/32 models on all eight tasks.
-Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
-Reference Methods
-Pre-trained 63.2 59.8 60.7 46.0 31.6 32.5 48.2 43.9 48.2
-Fine-tuned (STL) 75.0 78.3 95.2 99.0 97.3 98.9 99.6 79.7 90.3
-Multi-Task Model Fusion Methods
-Weight Averaging [57] 65.4 62.6 70.8 76.9 64.5 54.9 86.3 50.9 66.5
-Fisher Merging [37] 66.7 64.0 72.2 91.6 69.0 64.3 83.5 53.7 70.6
-RegMean [28] 67.8 68.9 82.5 94.4 90.6 79.2 97.6 63.2 80.5
-Task Arithmetic [25] 57.1 55.7 64.9 76.7 77.9 68.5 96.1 47.2 68.0
-Ties-Merging [62] 67.1 64.2 74.1 76.8 77.7 69.4 94.1 54.0 72.2
-task-wise AdaMerging [63] 58.6 56.9 69.8 82.4 70.3 58.9 97.2 55.3 68.7
-layer-wise AdaMerging [63] 67.9 71.3 83.5 92.7 87.4 92.9 98.2 67.0 82.6
-WEMoE (Model Mixing) [51] 73.7 76.8 93.4 98.2 96.8 98.2 99.6 76.6 89.2
-Table 4: Experimental results of merging single-task Resnet50 models on three NYUv2 tasks.
-Method Segmentation Depth Estimation Normal
-mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Mean $\downarrow$
-Single-Task Learning
-Segmentation 52.0 73.8 242.8 88.7 82.8
-Depth Estimation 2.3 6.2 42.5 17.7 82.8
-Normal 2.0 4.9 264.0 98.1 24.7
-Multi-Task Model Fusion Methods
-Weight Averaging [57] 39.0 67.0 55.1 22.7 30.4
-Task Arithmetic ($\lambda=0.3$[25] 33.6 63.3 56.3 23.2 31.3
-Ties-Merging ($\lambda=0.3$[62] 36.3 61.7 60.5 24.5 33.1
-Table 5: Generalization results on two unseen tasks when merging ViT-B/32 models on six tasks.
-Method Seen Tasks Unseen Tasks
-SUN397 Cars RESISC45 DTD SVHN GTSRB Avg. MNIST EuroSAT Avg.
-Pre-trained 63.2 59.9 60.6 43.9 23.5 30.4 46.9 47.6 45.6 46.6
-Fisher Merging [37] 65.5 67.2 78.2 57.6 84.2 75.9 71.4 71.8 49.4 60.6
-RegMean [28] 68.7 70.0 86.5 65.9 93.9 86.7 78.6 82.2 49.3 65.7
-Task Arithmetic [25] 64.3 63.0 73.2 54.9 84.7 79.5 69.9 75.5 42.6 59.1
-Ties-Merging [62] 68.3 65.5 76.9 54.9 75.4 72.0 68.9 73.1 47.3 60.2
-AdaMerging [63] 68.4 71.9 87.9 69.1 92.2 93.8 80.5 77.7 47.3 62.5
-WEMoE [51] 75.4 77.5 94.3 77.0 96.8 98.7 86.6 78.3 44.0 61.1
-Table 6: Ablations of the test data distribution on ViT-B/32 (for all methods, $\lambda=0.3$).
-Method Cars EuroSAT RESISC45 GTSRB Avg. Cars EuroSAT RESISC45 GTSRB Avg.
-Clean Test Set Corrupted Test Set (Motion Blur)
-Fisher Merging [37] 66.0 92.7 83.7 78.7 80.3 60.7 57.6 81.7 78.4 69.6
-RegMean [28] 72.1 97.5 88.9 93.9 88.1 70.0 71.3 87.5 86.8 78.9
-Task Arithmetic [25] 64.6 91.8 80.2 74.8 77.9 62.4 59.2 78.5 63.3 65.9
-Ties-Merging [62] 65.2 83.3 78.1 67.4 73.5 64.4 53.9 76.4 57.1 62.9
-AdaMerging [63] 75.2 94.3 87.6 96.7 88.5 72.4 72.7 85.3 94.3 81.2
-WEMoE [51] 77.4 98.9 94.4 99.0 92.4 76.5 74.2 93.7 97.4 85.5
-Corrupted Test Set (Impluse Noise) Corrupted Test Set (Gaussian Noise)
-Fisher Merging [37] 61.5 50.0 74.7 52.6 59.7 61.6 48.1 76.0 51.3 59.3
-RegMean [28] 66.9 51.0 80.6 68.7 66.8 69.4 41.8 84.0 67.7 65.7
-Task Arithmetic [25] 59.8 53.3 72.3 45.0 57.6 61.5 52.5 75.0 50.1 59.8
-Ties-Merging [62] 60.2 45.6 69.8 38.3 53.5 61.8 47.3 73.1 42.3 56.1
-AdaMerging [63] 69.2 40.0 79.6 83.3 68.0 70.0 53.3 82.1 80.0 71.4
-WEMoE [51] 75.1 9.7 91.5 91.8 67.0 76.5 9.6 92.7 88.7 66.8
-Corrupted Test Set (Pixelate) Corrupted Test Set (Spatter)
-Fisher Merging [37] 2.2 34.0 17.0 63.2 29.1 61.4 64.2 74.6 47.3 61.9
-RegMean [28] 2.3 38.3 18.2 89.4 37.0 67.7 60.0 81.3 81.9 72.7
-Task Arithmetic [25] 2.3 33.2 19.1 65.6 30.0 61.0 62.5 72.8 57.0 63.3
-Ties-Merging [62] 3.3 31.8 18.0 58.5 27.9 61.3 52.9 70.3 48.1 58.2
-AdaMerging [63] 1.3 52.9 21.0 91.0 41.5 68.4 55.9 78.3 92.3 73.7
-WEMoE [51] 0.5 11.6 2.3 97.5 28.0 75.1 9.7 91.4 96.3 68.1
-Corrupted Test Set (Contrast) Corrupted Test Set (JPEG Compression)
-Fisher Merging [37] 63.8 58.4 75.5 70.4 67.0 66.3 67.6 82.6 58.9 68.8
-RegMean [28] 69.6 64.8 84.4 90.0 77.2 71.5 72.6 88.7 82.2 78.7
-Task Arithmetic [25] 62.3 55.7 75.3 70.8 66.0 63.9 66.1 80.1 61.0 67.8
-Ties-Merging [62] 64.2 52.4 74.8 63.5 63.7 65.0 59.5 77.9 53.2 63.9
-AdaMerging [63] 73.1 67.4 83.0 96.2 79.9 72.9 70.7 86.3 90.6 80.1
-WEMoE [51] 77.2 34.7 93.1 98.4 75.9 77.3 61.0 94.1 95.7 82.0
-In this evaluation, we begin by comparing multi-task model fusion algorithms under several settings:
+Model mixing methods can be implemented through layer recombinations (hu2023fedmr; jiang2024evomerge), model stitching (lenc2015understanding; moschella2022relative), or upscale to create a Mixture of Experts (MoE)-based sparse model (komatsuzaki2022sparse; ye2023taskexpert; tang2024merging; lu2024twin; dai2024deepseekmoe; zhao2024loraretriever; ostapenko2024towards; tang2024smile; yadav2024survey).
+A.2 Related Benchmarks and Toolkits
+Although several model fusion methods have been proposed, comprehensive benchmarks and unified toolkits are still lacking in this field.
+A recent notable work, MergeKit (goddard2024arcee), provides a collection of model fusion techniques specifically designed for merging large language models (LLMs), focusing primarily on methods for Transformer-based architectures.
+While MergeKit excels in providing practical tools for LLM merging, its scope is inherently limited to a single domain (natural language processing) and a specific model family (Transformers).
+Similarly, MergeBench (he2025mergebench) is targeted exclusively at evaluating LLM merging techniques, providing benchmarks only within the language modeling domain.
+In contrast, FusionBench addresses several critical gaps in the current landscape, making FusionBench a more generalized and versatile platform for systematically assessing the performance of model fusion approaches across various domains and architectures: (1) Broader Domain Coverage: Unlike existing tools that focus solely on NLP, FusionBench spans both computer vision and natural language processing domains, enabling cross-domain evaluation and comparison.
+(2) Comprehensive Taxonomy: While MergeKit and MergeBench concentrate primarily on model merging methods, FusionBench encompasses a complete taxonomy including model ensemble, model merging, and model mixing methods, providing a unified framework for all major fusion paradigms.
+(3) Research-Oriented Evaluation: FusionBench is designed as a research platform that includes diverse fine-tuned models, standardized evaluation protocols, and comparative analysis utilities across different architectures (CNNs, Vision Transformers, CLIP models, etc.) and tasks.
+(4) Algorithmic Diversity: Beyond basic merging techniques, FusionBench incorporates advanced methods such as subspace-based fusion, gradient alignment, and mixture-of-experts approaches.
+Appendix B Project Structure
+To provide a clear overview of FusionBench’s organization, we present its comprehensive project structure in Listing 1.
+This structure reflects the modular design philosophy of our benchmark, with distinct directories for configuration management, documentation, example implementations, core functionality, and testing.
+1|-- config/ # Configuration files managed by Hydra
+2| |-- method/ # Configs for implemented algorithms
+3| |-- modelpool/ # Configs for model pool
+4| |-- taskpool/ # Configs for task pool
+5| |-- model/ # Configs for models
+6| |-- dataset/ # Configs for datasets
+7|-- docs/ # Documentation (mkdocs)
+8|-- examples/ # Example scripts and notebooks
+9| |-- {method_name}/ # Method-specific examples
+10|-- fusion_bench/ # Main package
+11| |-- method/ # Fusion algorithm implementations
+12| | |-- {method_name}/ # Method-specific code
+13| |-- modelpool/ # Model pool implementations
+14| |-- taskpool/ # Task pool implementations
+15| |-- ... # Models, datasets, and other utilities
+16|-- tests/ # Unit tests
+Listing 1: Project structure of FusionBench.
+Figure 3: Flowchart of FusionBench.
+Appendix C FusionBench Architecture and Implementation Guide
+Figure 3 illustrates the steps to fuse a model in FusionBench across three main stages: creating a model pool, instantiating and running an algorithm, and (optional) model evaluation.
+Pseudocode Example.
+The pseudocode for the pipeline is shown in Algorithm 1.
+Algorithm 1 Pseudocode for FusionBench main pipeline
+1:Load and parse configuration from YAML files
+2:Instantiate fusion_algorithm, model_pool, and (optionally) task_pool
+[…]
+4:Save merged_model if save path is specified
+5:if task pool is specified then
+6:  $\textit{report}\leftarrow\textit{task\_pool.evaluate}(\textit{merged\_model})$
+7:  Save report if save path is specified
+8:end if
+This design allows users to flexibly experiment with different algorithms, models, and tasks by simply modifying the configuration, without changing the core program logic.
+C.1 Implementing Customized Algorithm in FusionBench
+Here, we provide an example of implementing a customized algorithm in Listing 2.
+The derived algorithm class should inherit from the “BaseAlgorithm” class and implement the “run” method.
+The “run” method takes a “BaseModelPool” object as input, which is responsible for managing the models and dataset to be loaded.
+The “run” method should be implemented to fuse the models and return the fused model.
+1from fusion_bench import BaseAlgorithm, BaseModelPool
+2
+3class DerivedAlgorithm(BaseAlgorithm):
+4 """
+5 An example of a derived algorithm class.
+6 """
+7
+8 # _config_mapping maps the attribution to the corresponding key in the configuration file.
+9 # this is optional and can be used to serialize the object to a configuration file.
+10 # ‘self.config.hyperparam_1‘ will be mapped to the attribute ‘hyperparam_attr_1‘.
+11 _config_mapping = BaseModelFusionAlgorithm._config_mapping | {
+12 "hyperparam_attr_1": "hyperparam_1",
+13 "hyperparam_attr_2": "hyperparam_2",
+14 }
+15
+16 def __init__(self, hyperparam_1, hyperparam_2, **kwargs):
+17 super().__init__(**kwargs)
+18 self.hyperparam_attr_1 = hyperparam_1
+19 self.hyperparam_attr_2 = hyperparam_2
+20
+21 def run(self, modelpool: BaseModelPool):
+22 # modelpool is an object that responsible for managing the models and dataset to be loaded.
+23 # implement the fusion algorithm here.
+24 raise NotImplementedError(
+25 "DerivedAlgorithm.run() is not implemented."
+26 )
+Listing 2: Implementing new algorithm in FusionBench.
+As an illustration, we can implement a deep model fusion algorithm that merges models through parameter averaging (wortsman2022model).
+The following code snippet demonstrates this approach (Listing 3):
+1class SimpleAverageAlgorithm(
+2 BaseAlgorithm,
+3 SimpleProfilerMixin,
+4):
+5 @torch.no_grad()
+6 def run(self, modelpool: Union[BaseModelPool, Dict[str, $n$.Module]]):
+7 """
+8 Fuse the models in the given model pool using simple averaging.
+9
+10 This method iterates over the names of the models in the model pool, loads each model, and appends it to a list.
+11 It then returns the simple average of the models in the list.
+12
+13 Args:
+14 modelpool: The pool of models to fuse.
+15
+16 Returns:
+17 The fused model obtained by simple averaging.
+18 """
+19 if isinstance(modelpool, dict):
+20 modelpool = BaseModelPool(modelpool)
+21
+22 log.info(
+23 f"Fusing models using simple average on {len(modelpool.model_names)} models."
+24 f"models: {modelpool.model_names}"
+25 )
+26 sd: Optional[StateDictType] = None
+27 forward_model = None
+28 merged_model_names = []
+29
+30 for model_name in modelpool.model_names:
+31 with self.profile("load model"):
+32 model = modelpool.load_model(model_name)
+33 merged_model_names.append(model_name)
+34 print(f"load model of type: {type(model).__name__}")
+35 with self.profile("merge weights"):
+36 if sd is None:
+37 # Initialize the state dictionary with the first model’s state dictionary
+38 sd = model.state_dict(keep_vars=True)
+39 forward_model = model
+40 else:
+41 # Add the current model’s state dictionary to the accumulated state dictionary
+42 sd = state_dict_add(sd, model.state_dict(keep_vars=True))
+43 with self.profile("merge weights"):
+44 # Divide the accumulated state dictionary by the number of models to get the average
+45 sd = state_dict_div(sd, len(modelpool.model_names))
+46
+47 forward_model.load_state_dict(sd)
+48 # print profile report and log the merged models
+49 self.print_profile_summary()
+50 log.info(f"merged {len(merged_model_names)} models:")
+51 for model_name in merged_model_names:
+52 log.info(f" - {model_name}")
+53 return forward_model
+Listing 3: Implementation of simple average algorithm in FusionBench.
+C.2 Customizing Model Pools in FusionBench
+The Model Pool Module serves as an abstraction layer that manages model loading, saving, and dataset handling during fusion operations.
+To extend FusionBench to new model architectures, users need to implement a custom model pool class that inherits from BaseModelPool and defines architecture-specific operations.
+This design principle ensures that fusion algorithms remain architecture-agnostic and can be applied universally across different model types without modification.
+Extensibility to Modern Architectures.
+FusionBench’s modular architecture facilitates straightforward integration of emerging model architectures.
+The key advantage of this design is that once a model pool class is implemented for a particular architecture, all existing fusion algorithms in the benchmark become immediately applicable to that architecture.
+This extensibility has enabled us to expand beyond our initial focus on established models (CLIP-ViT, ResNet-50, GPT-2, and Flan-T5) to encompass state-of-the-art architectures across multiple domains.
+For computer vision tasks, we have integrated modern architectures including ConvNeXt (liu2022convnet) and DINOv2 (oquab2024dinov), both of which represent significant advances in efficient visual representation learning.
+These models can be loaded through dedicated model pool classes (ConvNeXtForImageClassificationPool and DINOv2ForImageCla-
+ssificationPool
+) with comprehensive online documentation available on our project website.
+For natural language processing tasks, our CausalLMPool class provides broad compatibility with contemporary large language model architectures.
+The CausalLMPool leverages the HuggingFace Transformers library, allowing seamless integration of new architectures as they are released.
+Appendix D Fine-Tuned Model Collections in FusionBench
+This section outlines the comprehensive model collections provided in FusionBench, including their experimental configurations and performance evaluation results across various single-task specializations.
   1. 1.
-    Image Classification: We use CLIP models from the HuggingFace Library [24]. The results are in Table 3 and Figure 4 for CLIP-ViT-B/32, and in Table 14 for CLIP-ViT-L/14.
-  2. 2.
-    Scene Understanding: Using ResNet-50 models on the NYUv2 dataset for segmentation, depth estimation, and normal estimation tasks. Results are in Table 4.
-  3. 3.
-    Text Classification: Results for GPT-2 models on seven tasks are shown in Table 12.
-  4. 4.
-    Text-to-Text Generation: For LoRA fine-tuned Flan-T5-base and Flan-T5-large models, after merging and unloading the LoRA adapters, results are in Tables 13 and 15.
-In these tables, we compare the performance of different multi-task model fusion algorithms across various tasks.
-Pre-trained models’ performance, fine-tuned models’ performance, and traditional multi-task learning (MTL) methods are provided for reference.
-In Appendix A, we provide a detailed description of these fine-tuned single-task models.
-We have the following key observations: (1) Multi-task model fusion algorithms generally outperform pre-trained models.
-This indicates that multi-task model fusion can effectively transfer knowledge from multiple single-task models to improve performance on each task.
-(2) Layer-wise AdaMerging and Weight-Ensembling MoE achieve the best overall performance among the multi-task model fusion methods.
-These methods demonstrate the effectiveness of adaptive layer-wise merging and mixture-of-experts approaches in multi-task model fusion.
-(3) The performance gap between multi-task model fusion and fine-tuned models (STL) is larger for CLIP-ViT-B/32 compared to CLIP-ViT-L/14.
-This suggests that multi-task model fusion may be more beneficial for smaller models, as they have more room for improvement through knowledge transfer.
-(4) Traditional MTL outperforms most multi-task model fusion methods, which indicates that traditional MTL is still a strong baseline for multi-task learning, and there is room for improvement in multi-task model fusion algorithms.
-4.3 Generalization and Robustness Evaluation
-To further assess the generalization and robustness of multi-task model fusion algorithms, we conduct experiments on unseen tasks and corrupted test sets (or out-of-distribution test sets).
-(1) Tables 5 and 16 present the generalization performance of various multi-task model fusion algorithms when merging CLIP-ViT-B/32 models trained on six seen tasks and evaluating their performance on two unseen tasks.
-This analysis helps us understand how well the fused models can adapt to new tasks that were not encountered during the training and model fusion process.
-Additional details and discussions regarding the generalization experiments can be found in Appendix C.
-(2) Furthermore, in Table 6, we investigate the robustness of multi-task model fusion algorithms by evaluating their performance on corrupted test sets.
-These corrupted test sets are designed to simulate real-world scenarios where the input data may be noisy or corrupted.
-We have the following key observations: (1) The performance of all multi-task model fusion methods on unseen tasks is generally lower than their performance on seen tasks.
-This is expected, as the models being fused are not explicitly trained on the unseen tasks.
-(2) There is a negative transfer observed on the RESISC45 dataset in the generalization experiments.
-The performance of all multi-task model fusion methods on RESISC45 is lower than the pre-trained model, indicating that the knowledge transferred from the seen tasks may not be beneficial or even harmful to this specific unseen task.
-(3) The performance of all methods drops significantly on certain types of corruptions, such as pixelation and impulse noise.
-This highlights the challenge of maintaining robustness under severe distribution shifts and the need for further research in this direction.
-(4) When the test distribution is corrupted, adaptive methods may overfit to certain tasks, leading to a decrease in overall performance.
-This suggests that adaptive methods may need to be further regularized to improve generalization and robustness.
-5 Conclusions, Future Plans and Societal Impacts
-Conclusions.
-We’ve developed a flexible and modular codebase, which serves as the foundation for FusionBench.
-Our benchmark provides a comprehensive evaluation framework for assessing the performance of multi-task model fusion algorithms.
-This innovative and comprehensive framework underscores the advantages of a scalable and extendable architecture, thereby simplifying the creation of deep model fusion algorithms.
-We also organize and provide a collection of datasets and models, which can be utilized to ensure a fair comparison.
-Last, FusionBench comes with extensive documentation and a series of tutorials, making it user-friendly for beginners and interested researchers.
-We hope that the community will leverage this benchmark to develop and evaluate new fusion algorithms and to further the popularity of deep model fusion in the machine learning community.
-Limitations and Future Plans.
-To date, FusionBench primarily focuses on the evaluation of deep model fusion algorithms for multi-task learning.
-Despite having implemented numerous fusion algorithms, including those that don’t primarily focus on multi-task learning, we have not yet to investigate the evaluation for these methods.
-In the future, we plan to extend the benchmark to provide a more comprehensive evaluation framework for them.
-What’s more, we plan to extend the benchmark by incorporating additional datasets and applications, such as human preference alignment, multi-modal fusion, and reinforcement learning tasks.
-Societal Impacts and Ethical Issue.
-Our benchmark has the potential to accelerate the development of deep model fusion algorithms.
-This can play a crucial role in speeding up the development of machine learning models.
-By optimizing and combining models more efficiently, we can substantially reduce the carbon footprint associated with training these models, thereby making a positive contribution to environmental sustainability.
-Given these substantial benefits, the potential negative societal impacts of our work are minimal.
-Appendix A Evaluation of Fine-Tuned Single-Task Models
-In this section, we describe the experimental setup for fine-tuning the models and present the performance results of the fine-tuned single-task models.
-  1. 1.
-    CLIP-ViT-B/32 Models: The CLIP models are fine-tuned on eight image classification tasks: SUN397, Cars, RESISC45, EuroSAT, SVHN, GTSRB, MNIST, and DTD.
-The Adam Optimizer is employed with a fixed learning rate of 1e-5 for a total of 4000 training steps with the batch size of 32.
-Only the vision encoder is fine-tuned to maintain the model’s open-vocabulary characteristic.
-The performance of fine-tuned CLIP-ViT-B/32 and CLIP-ViT-L/14 models on the eight image classification tasks is shown in Tables 7 and 8, respectively.
-In Figure 5, we visualize the cosine similarity matrices of task vectors for CLIP-ViT-B/32 and CLIP-ViT-L/14 models.
-We note that the task vectors for models from various tasks are nearly orthogonal.
-This suggests that the knowledge specific to each task resides in distinct directions or subspaces.
-This finding motivates the exploration of locating subspaces in which the knowledge of different tasks can be merged effectively, as discussed in Tang et al.
-[50].
+    CLIP-ViT Models: In FusionBench, we provide comprehensive CLIP-ViT model collections fine-tuned on up to 20 image classification tasks.
+The most commonly used eight tasks are SUN397 (xiao2010sun), Cars (krause20133d), RESISC45 (cheng2017remote), EuroSAT (helber2018introducing), SVHN (netzer2011reading), GTSRB (stallkamp2012man), MNIST (lecunGradientbasedLearningApplied1998), and DTD (cimpoi2014describing).
+Additionally, we extend our collection to include 12 more tasks: Oxford Flowers102 (nilsback2008automated), PatchCamelyon (veeling2018rotation), FER2013 (goodfellow2013challenges), Oxford-IIIT Pet (parkhi2012cats), STL-10 (coates2011analysis), CIFAR-100 (krizhevsky2009learning), CIFAR-10, Food-101 (bossard2014food), Fashion-MNIST (xiao2017fashion), EMNIST Letters (cohen2017emnist), Kuzushiji-MNIST (clanuwat2018deep), and Rendered-SST2 (socher2013recursive), covering diverse domains from natural images to histopathologic data and facial expressions.
+We use the Adam Optimizer with a fixed learning rate of 1e-5 over 4000 training steps (batch size=32).
+Only the vision encoder is fine-tuned while the text encoder remains fixed to preserve the open-vocabulary property of the model.
+All fine-tuned models are publicly available on the HuggingFace Model Hub  [^1] https://huggingface.co/ and Modelscope  [^2] https://www.modelscope.cn/home for easy access and reproducibility.
+In Section D.1, we present the performance of fine-tuned single-task CLIP-ViT-B/32 and CLIP-ViT-L/14 models on the eight most commonly used eight image classification tasks.
   2. 2.
     ResNet-50 Models: We fine-tune ResNet-50 models on three scene understanding tasks: segmentation, depth estimation, and normal estimation using the NYUv2 dataset, each with a learning rate of 1e-4 for 40 epochs, the learning rate is reduced by a factor of 0.5 every 10 epochs.
-The performance of fine-tuned single-task ResNet-50 models on the NYUv2 dataset is shown in Table 4.
+The performance of fine-tuned single-task ResNet-50 models on the NYUv2 dataset is shown in Section D.2.
   3. 3.
-    GPT-2 Models: GPT2 model fine-tuned on tasks from GLUE benchmark, using a constant learning rate of 5e-5 for 3 epochs.
-The performance of fine-tuned single-task GPT-2 models on the seven text classification tasks is shown in Table 9.
+    GPT-2 Models: We fine-tune GPT-2 models on seven text classification tasks from the GLUE benchmark (wang2018glue) to evaluate their effectiveness in natural language understanding scenarios.
+The fine-tuning process employs the Adam optimizer with a constant learning rate of 5e-5 over 3 epochs, following standard practices for transformer-based language models.
+These tasks include CoLA, MNLI, MRPC, QNLI, QQP, RTE, and SST-2 .
+The performance evaluation results of these fine-tuned single-task GPT-2 models are presented in Section D.3.
   4. 4.
-    Flan-T5 Models: In this work, we fine-tune Flan-T5-base and Flan-T5-large models on eight text-to-text generation tasks from the GLUE benchmark.
-The results of LoRA fine-tuned Flan-T5-base and Flan-T5-large models are shown in Tables 10 and 11, respectively.
+    Flan-T5 Models: We fine-tune both Flan-T5-base and Flan-T5-large models on eight text-to-text generation tasks from the GLUE benchmark.
+For reproducibility and transparency, we provide the detailed prompt templates used for each task in Section F, ensuring that researchers can replicate our experimental setup and build upon our work.
+The performance evaluation results of these LoRA fine-tuned Flan-T5-base and Flan-T5-large models across all eight tasks are presented in Section D.4.
+Based on the performance metrics detailed in Section D.1, D.2, D.3 and D.4, we observe that the fine-tuned models demonstrate high accuracy on specific tasks.
+This observation holds true across various model architectures and task domains, indicating the effectiveness of the fine-tuning process in adapting pre-trained models to excel in particular applications.
+Furthermore, fine-tuning a model on one task can lead to both positive or negative transfer effects on other tasks.
+Positive transfer occurs when the knowledge gained from the target task enhances the model’s performance on another task, while negative transfer arises when the fine-tuning process on the target task hinders the model’s ability to perform on others.
+D.1 CLIP-ViT Models
 Table 2: Performance of fine-tuned single-task CLIP-ViT-B/32 models on the eight image classification tasks.
@@
 SVHN 40.5 36.3 18.9 9.8 97.3 27.3 81.8 23.2
-GRSRB 36.9 33.0 20.6 21.3 41.2 98.9 30.9 23.9
+GTSRB 36.9 33.0 20.6 21.3 41.2 98.9 30.9 23.9
 MNIST 50.3 40.0 31.3 17.7 50.1 19.3 99.6 30.7
@@
 SVHN 66.5 69.0 54.0 19.7 97.9 48.7 92.2 50.1
-GRSRB 63.4 64.8 38.7 19.6 71.0 99.2 75.1 45.8
+GTSRB 63.4 64.8 38.7 19.6 71.0 99.2 75.1 45.8
 MNIST 56.1 49.8 53.5 26.6 48.2 33.1 99.8 47.1
@@
 (b) CLIP-ViT-L/14 models.
-Figure 5: Cosine similarity matrices of task vectors for CLIP-ViT-B/32 and CLIP-ViT-L/14 models.
-Table 9: Performance of fine-tuned single-task GPT-2 models on the seven text classification tasks.
+Figure 4: Cosine similarity of task vectors for CLIP-ViT-B/32 and CLIP-ViT-L/14 models.
+The performance of fine-tuned CLIP-ViT-B/32 and CLIP-ViT-L/14 models on the eight most commonly used image classification tasks is shown in Tables 2 and 3, respectively.
+In Figure 4, we visualize the cosine similarity matrices of task vectors for CLIP-ViT-B/32 and CLIP-ViT-L/14 models.
+Note that while we present results for the eight core tasks in this paper for conciseness, comprehensive evaluation results for all 20 tasks across CLIP-ViT-B/32, CLIP-ViT-B/16, and CLIP-ViT-L/14 models are available in our online documentation.
+Following (ilharco2022editing), the task vector is defined as the difference between a fine-tuned model’s parameters and the original pre-trained model’s parameters: $\tau=\theta_{ft}-\theta_{pre}$, where $\theta_{ft}$ represents the fine-tuned weights and $\theta_{pre}$ represents the pre-trained weights.
+We note that the task vectors for models from various tasks are nearly orthogonal.
+This suggests that the knowledge specific to each task resides in distinct directions or subspaces.
+This finding motivates the exploration of locating subspaces in which the knowledge of different tasks can be merged effectively, as discussed in tang2023concrete; wang2024localizing.
+D.2 ResNet-50 Models
+Table 4: Single-task performance of fine-tuned ResNet-50 models on NYUv2 dataset.
+Method SEGMENTATION DEPTH ESTIMATION NORMAL
+mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Mean $\downarrow$
+Segmentation 52.0 73.8 242.8 88.7 82.8
+Depth Estimation 2.3 6.2 42.5 17.7 82.8
+Normal 2.0 4.9 264.0 98.1 24.7
+We fine-tune ResNet-50 models on three distinct scene understanding tasks using the NYUv2 dataset: semantic segmentation, depth estimation, and surface normal estimation.
+Each model is trained with a learning rate of 1e-4 for 40 epochs, with the learning rate reduced by a factor of 0.5 every 10 epochs.
+Table 4 presents the single-task performance of these specialized models, where each row represents a model trained exclusively on one task and evaluated across all three tasks using task-specific metrics.
+The results demonstrate clear task specialization: the segmentation model achieves optimal performance on segmentation metrics (52.0 mIoU, 73.8 Pixel Accuracy), the depth estimation model excels at depth-related metrics (42.5 Absolute Error, 17.7 Relative Error), and the normal estimation model performs best on surface normal prediction (24.7 Mean angular error).
+Notably, models trained on one task show severely poor performance on other tasks, highlighting the challenge of multi-task learning in computer vision applications.
+D.3 GPT-2 Models
+Table 5: Performance of fine-tuned GPT-2 models on the seven text classification tasks.
 Model CoLA MNLI MRPC QNLI QQP RTE SST-2
@@
 SST-2 51.8 32.9 40.2 49.8 56.8 44.4 91.2
+We evaluate GPT-2 models fine-tuned on seven distinct text classification tasks from the GLUE benchmark, as detailed in the enumeration above.
+Table 5 presents the cross-task performance evaluation, where each row represents a model fine-tuned on a specific task (indicated by the row label) and columns show its performance across all seven tasks.
+The results demonstrate clear task specialization, with each fine-tuned model achieving its highest performance on its target task (diagonal elements).
+D.4 Flan-T5 Models
 Table 6: Performance of LoRA fine-tuned Flan-T5-Base models on the eight text-to-text generation tasks from the GLUE benchmark.
@@
 STSB 73.4 39.3 82.1 92.6 86.1 83.4 94.0 90.9
-Based on the performance metrics detailed in these tables, we observe that the fine-tuned models demonstrate high accuracy on specific tasks.
-This observation holds true across various model architectures and task domains, indicating the effectiveness of the fine-tuning process in adapting pre-trained models to excel in particular applications.
-What’s more, fine-tuning a model on one task can lead to both positive and negative transfer effects on other tasks.
-Positive transfer occurs when the knowledge gained from fine-tuning on one task enhances the model’s performance on another task, while negative transfer arises when the fine-tuning process on one task hinders the model’s ability to perform well on other tasks.
-Appendix B Multi-Task Model Fusion
+We evaluate Flan-T5-base and Flan-T5-large models fine-tuned using LoRA on eight text-to-text generation tasks from the GLUE benchmark (wang2018glue).
+Tables 6 and 7 present the cross-task performance evaluation for both model sizes, where each row represents a model fine-tuned on a specific task and columns show performance across all eight tasks.
+The results demonstrate that each specialized model achieves its best performance on its target task (diagonal elements), with the Flan-T5-large models generally outperforming their base counterparts across all tasks.
+Notably, when evaluating fine-tuned models on non-target tasks, we observe both positive and negative transfer effects.
+For instance, in Flan-T5-base models, the MNLI-specialized model shows positive transfer to QNLI (89.3% vs 88.% pre-trained baseline), while the STSB-specialized model exhibits negative transfer to MNLI (1.7% vs 56.5% pre-trained baseline).
+Appendix E Experimental Evaluation and Benchmarking Results
+In this section, we evaluate the performance of multi-task deep model fusion algorithms on a variety of tasks, as well as analyze the generalization and robustness of these algorithms.
+We also provide an ablation study to investigate the impact of hyperparameter selection.
+The majority of experiments presented in this section can be executed using a single NVIDIA RTX 3090 GPU with 24GB memory, ensuring cost-effective reproducibility.
+E.1 Experimental Setup
+In this section, we conduct a series of multi-task deep model fusion experiments on image classification tasks, text classification tasks, and text-to-text generation tasks to evaluate the performance of multi-task deep model fusion algorithms.
+These tasks are chosen to cover a wide range of NLP and CV tasks, as described in Section E.3.
+E.2 Computational Cost Analysis
+While functional requirements listed in Table 1 differ across methods, computational resource implications also vary significantly.
+For instance, ensemble methods incur high inference costs since multiple models must be stored and executed simultaneously.
+Model merging methods are generally cost-efficient during inference, as the merged model retains the footprint of a single model, though some approaches require additional data-dependent computations during merging.
+Mixing methods typically expand the model size and necessitate partial re-training, leading to substantial computational overhead.
+Practical Implications: For resource-constrained environments, parameter-free methods like parameter average and task arithmetic offer the best cost-performance trade-off.
+When additional data is available, alignment-based methods such as RegMean (jin2022dataless) and RegMean++ (nguyen2025regmean++) provide strong performance with moderate computational overhead.
+For applications requiring maximum performance and having sufficient computational resources or access to the test data distribution, test-time adaptive methods like AdaMerging (yang2023adamerging) or model mixing approaches like WE-MoE (tang2024merging) are recommended despite their higher computational costs.
+Besides, the LazyStateDict utility in FusionBench significantly reduces memory requirements for all methods by enabling on-demand parameter loading, making large-scale model fusion practical.
+E.3 Multi-Task Deep Model Fusion
+We start our evaluation by comparing multi-task model fusion algorithms across multiple experimental settings, including both computer vision and natural language processing tasks:
+  1. 1.
+    Open-Vocabulary Image Classification: Using the CLIP-ViT models fine-tuned on the eight most commonly used tasks. Results for CLIP-ViT-B/32 and CLIP-ViT-L/14 models are presented in Section E.3.1.
+  2. 2.
+    Scene Understanding: Using ResNet-50 models on the NYUv2 dataset for segmentation, depth estimation, and normal estimation tasks. Detailed results are provided in Section E.3.2.
+  3. 3.
+    Text Classification: Detailed results for GPT-2 models on seven text classification tasks are presented in Section E.3.3.
+  4. 4.
+    Text-to-Text Generation: For Flan-T5-base and Flan-T5-large models fine-tuned with LoRA, the results are presented in Section E.3.4.
+In Section E.3.1, E.3.2, E.3.3 and E.3.4, we compare the performance of different multi-task model fusion algorithms across various tasks.
+Pre-trained models’ performance, fine-tuned models’ performance, and traditional multi-task learning (MTL) methods are provided for reference.
+In Appendix D, we provide a detailed description of these fine-tuned single-task models.
+We have the following key observations: (1) Multi-task model fusion usually outperforms pre-trained models, showing it can transfer knowledge from multiple single-task models to enhance performance.
+Pre-trained models lack task-specific knowledge as they are not fine-tuned for downstream tasks.
+(2) Adaptive method (AdaMerging) and MoE-based method perform best in multi-task model fusion, showing the effectiveness of adaptive merging and mixture-of-experts approaches.
+(3) The performance gap between multi-task model fusion and single-task fine-tuned models (STL) is larger for CLIP-ViT-B/32 compared to CLIP-ViT-L/14.
+This suggests that multi-task model fusion may be more beneficial for smaller models, as they have more room for improvement through knowledge transfer.
+(4) Traditional MTL outperforms most multi-task model fusion methods, which indicates that traditional MTL is still a strong baseline for multi-task learning, and there is room for improvement in multi-task model fusion algorithms.
+E.3.1 Open-Vocabulary Image Classification Using CLIP-ViT Models
+Figure 5: Radar charts comparing the performance of different deep model fusion methods across eight tasks using CLIP-ViT-B/32 and CLIP-ViT-L/14.
+Table 8: Multi-task performance when merging CLIP-ViT-B/32 models on all eight tasks.
+Here underlined method names indicate methods that require test-time adaptation training, while italicized method names indicate model mixing approaches.
+Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD AVG.
+Reference Methods
+Pre-trained 63.2 59.8 60.7 46.0 31.6 32.5 48.2 43.9 48.2
+Individual Fine-tuned 75.0 78.3 95.2 99.0 97.3 98.9 99.6 79.7 90.3
+Traditional MTL 72.3 76.6 92.2 97.9 95.5 97.7 99.3 77.7 88.6
+Multi-Task Model Fusion Methods
+Weight Averaging 65.4 62.6 70.8 76.9 64.5 54.9 86.3 50.9 66.5
+Fisher Merging 66.7 64.0 72.2 91.6 69.0 64.3 83.5 53.7 70.6
+RegMean 68.6 70.0 84.6 95.4 92.6 83.4 98.4 66.1 82.4
+RegMean++ 69.3 70.5 86.7 96.1 94.1 90.4 99.0 68.7 84.4
+Task Arithmetic 57.1 55.7 64.9 76.7 77.9 68.5 96.1 47.2 68.0
+Ties-Merging 67.1 64.2 74.1 76.8 77.7 69.4 94.1 54.0 72.2
+task-wise AdaMerging 58.6 56.9 69.8 82.4 70.3 58.9 97.2 55.3 68.7
+layer-wise AdaMerging 67.9 71.3 83.5 92.7 87.4 92.9 98.2 67.0 82.6
+WEMoE (Model Mixing) 73.7 76.8 93.4 98.2 96.8 98.2 99.6 76.6 89.2
+SMILE (Model Mixing) 73.6 77.8 92.0 98.3 96.9 98.1 99.6 78.1 89.3
+Table 9: Multi-task performance when merging CLIP-ViT-L/14 models on all eight tasks.
+Here underlined method names indicate methods that require test-time adaptation training, while italicized method names indicate model mixing approaches.
+Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
+Reference Methods
+Pre-trained 68.3 77.8 71.0 58.9 58.4 50.6 76.4 55.5 64.6
+Individual Fine-tuned 82.8 92.9 97.4 99.2 97.9 99.2 99.8 85.5 94.3
+Traditional MTL 79.0 89.3 94.5 98.4 96.4 98.1 99.4 83.7 92.4
+Multi-Task Model Fusion Methods
+Weight Averaging 72.5 81.5 82.2 90.0 81.6 74.0 96.6 61.8 80.0
+Fisher Merging 70.6 79.4 84.1 98.1 74.7 85.0 89.5 61.0 80.3
+RegMean 76.9 89.8 93.0 97.5 96.3 94.1 98.7 77.0 90.4
+RegMean++ 77.2 89.6 92.8 97.5 96.9 96.3 99.2 78.4 91.0
+Task Arithmetic 72.0 79.0 80.5 86.0 87.5 83.5 98.0 58.8 80.7
+Ties-Merging 74.7 83.3 86.4 91.3 89.7 85.2 97.8 63.9 84.0
+task-wise AdaMerging 75.8 80.1 77.2 83.6 68.4 93.5 93.1 69.0 80.1
+layer-wise AdaMerging 78.1 90.7 90.8 96.5 94.8 97.5 98.6 81.3 91.0
+WEMoE (Model Mixing) 81.5 92.3 96.5 98.8 97.6 99.4 99.6 84.5 93.8
+SMILE (Model Mixing) 81.9 92.3 95.5 99.1 98.0 98.9 99.7 83.6 93.6
+We evaluate model fusion methods on open-vocabulary image classification using CLIP-ViT models fine-tuned on eight diverse image classification tasks spanning natural scenes, satellite imagery, handwritten digits, and texture recognition.
+Tables 8 and 9 present comprehensive results for CLIP-ViT-B/32 and CLIP-ViT-L/14 models respectively, while Figure 5 provides radar chart visualizations highlighting the performance patterns across tasks.
+The results demonstrate several key insights: (1) Model mixing methods (such as WEMoE and SMILE) consistently achieve the best performance, approaching traditional multi-task learning results with averages of 89.2-89.3% for ViT-B/32 and 93.6-93.8% for ViT-L/14.
+(2) Layer-wise adaptive merging (AdaMerging) significantly outperforms task-wise variants, highlighting the importance of fine-grained parameter-level fusion strategies.
+(3) Simple parameter averaging and Fisher merging provide strong baselines but fall short of more sophisticated approaches like RegMean, which leverages data alignment techniques.
+(4) The performance gap between fusion methods and individual fine-tuned models is more pronounced for the smaller ViT-B/32 (90.3% vs 89.3%) compared to ViT-L/14 (94.3% vs 93.8%), suggesting that larger models may be more amenable to effective knowledge combination through fusion techniques.
+This phenomenon has also been discussed in the literature on large language model merging (yadav2024matters).
+E.3.2 Scene Understanding Using ResNet-50 Models
 Figure 6: Merging ResNet-50 models on three scene understanding tasks: segmentation, depth estimation, and normal estimation. Where the backbones are merged and the heads are kept separate.
-Table 12: Multi-task performance when merging GPT-2 models on seven text classification tasks.
+Table 10: Experimental results of merging single-task Resnet50 models on three NYUv2 tasks.
+Method SEGMENTATION DEPTH ESTIMATION NORMAL
+mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Mean $\downarrow$
+Single-Task Learning
+Segmentation 52.0 73.8 242.8 88.7 82.8
+Depth Estimation 2.3 6.2 42.5 17.7 82.8
+Normal 2.0 4.9 264.0 98.1 24.7
+Multi-Task Model Fusion Methods
+Weight Averaging 39.0 67.0 55.1 22.7 30.4
+Task Arithmetic ($\lambda=0.3$) 33.6 63.3 56.3 23.2 31.3
+Ties-Merging ($\lambda=0.3$) 36.3 61.7 60.5 24.5 33.1
+Figure 6 illustrates the multi-task model fusion problem in scene understanding, where three specialized ResNet-50 models trained individually on segmentation, depth estimation, and normal estimation tasks are merged at the backbone level while preserving task-specific heads.
+Table 10 presents the results.
+The results show that single-task ResNet50 models excel only at their own objectives but fail badly on other tasks, highlighting strong specialization without transferability.
+In contrast, model fusion produces balanced performance across tasks: for example, while fusion methods sacrifice some segmentation accuracy compared to the segmentation expert, they dramatically improve depth and normal performance relative to non-specialists.
+Table 11: Multi-task performance of GPT-2 models on seven text classification tasks.
 Method CoLA MNLI MRPC QNLI QQP RTE SST-2 Avg.
@@
 Fine-tuned (STL) 76.8 82.1 80.4 88.3 89.6 65.3 91.2 82.0
-Multi-Task Model Fusion Method
-Simple Average [57] 55.0 55.1 51.0 57.6 76.7 44.8 52.5 56.1
-Fisher Merging [37] 54.8 58.0 39.5 63.3 81.5 49.1 64.7 58.7
-RegMean [28] 61.7 70.4 65.4 69.7 78.8 56.0 79.7 68.8
-Task Arithmetic [25] 68.7 68.6 69.6 70.5 81.8 47.3 83.6 70.0
-Ties-Merging [62] 68.4 71.4 68.4 69.6 82.4 47.7 81.8 70.0
+Multi-Task Deep Model Fusion Method
+Simple Average 55.0 55.1 51.0 57.6 76.7 44.8 52.5 56.1
+Fisher Merging 54.8 58.0 39.5 63.3 81.5 49.1 64.7 58.7
+RegMean 61.7 70.4 65.4 69.7 78.8 56.0 79.7 68.8
+Task Arithmetic 68.7 68.6 69.6 70.5 81.8 47.3 83.6 70.0
+Ties-Merging 68.4 71.4 68.4 69.6 82.4 47.7 81.8 70.0
+E.3.3 Text Classification Using GPT-2 Models
+We evaluate multi-task model fusion methods on natural language understanding tasks using GPT-2 models fine-tuned on seven diverse text classification tasks from the GLUE benchmark (wang2018glue).
+Table 11 presents the results.
+E.3.4 Text-to-Text Generation Using Flan-T5 Models
 Table 12: Experimental results of merging Flan-T5-base (LoRA fine-tuned) models on all eight tasks.
+Here underlined method names indicate methods that require test-time adaptation training, while italicized method names indicate model mixing approaches.
 Method CoLA MNLI MRPC QNLI QQP RTE SST2 STSB Avg.
@@
 Multi-Task Model Fusion Methods
-Weight Averaging [57] 69.7 59.7 78.9 90.1 83.8 80.5 91.2 72.0 78.2
-Task Arithmetic [25] 68.8 55.2 78.7 89.8 83.7 79.1 91.5 72.4 77.4
-Ties-Merging [62] 68.3 56.3 79.4 89.8 83.7 79.4 91.6 71.2 77.5
-Layer-wise AdaMerging [63] 69.1 60.3 78.4 90.0 83.6 79.1 91.6 74.1 78.3
-Table 14: Multi-task performance when merging CLIP-ViT-L/14 models on all eight tasks.
-Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
-Reference Methods
-Pre-trained 68.3 77.8 71.0 58.9 58.4 50.6 76.4 55.5 64.6
-Fine-tuned (STL) 82.8 92.9 97.4 99.2 97.9 99.2 99.8 85.5 94.3
-Multi-Task Model Fusion Methods
-Weight Averaging [57] 72.5 81.5 82.2 90.0 81.6 74.0 96.6 61.8 80.0
-Fisher Merging [37] 70.6 79.4 84.1 98.1 74.7 85.0 89.5 61.0 80.3
-RegMean [28] 75.3 88.4 90.0 97.1 95.9 92.4 98.5 72.6 88.8
-Task Arithmetic [25] 72.0 79.0 80.5 86.0 87.5 83.5 98.0 58.8 80.7
-Ties-Merging [62] 74.7 83.3 86.4 91.3 89.7 85.2 97.8 63.9 84.0
-task-wise AdaMerging [63] 75.8 80.1 77.2 83.6 68.4 93.5 93.1 69.0 80.1
-layer-wise AdaMerging [63] 78.1 90.7 90.8 96.5 94.8 97.5 98.6 81.3 91.0
-WEMoE (Model Mixing) [51] 81.5 92.3 96.5 98.8 97.6 99.4 99.6 84.5 93.8
+Weight Averaging 69.7 59.7 78.9 90.1 83.8 80.5 91.2 72.0 78.2
+Task Arithmetic 68.8 55.2 78.7 89.8 83.7 79.1 91.5 72.4 77.4
+Ties-Merging 68.3 56.3 79.4 89.8 83.7 79.4 91.6 71.2 77.5
+Layer-wise AdaMerging 69.1 60.3 78.4 90.0 83.6 79.1 91.6 74.1 78.3
+SMILE (Model Mixing) 69.3 82.9 83.8 90.6 83.9 83.4 93.1 85.1 84.0
 Table 13: Experimental results of merging Flan-T5-large (LoRA fine-tuned) models on all eight tasks.
@@
 Multi-Task Model Fusion Methods
-Weight Averaging [57] 74.6 84.3 84.1 92.8 86.3 87.4 94.8 88.0 86.5
-Task Arithmetic [25] 76.9 85.4 85.3 93.9 85.8 88.1 95.2 87.8 87.3
-Ties-Merging [62] 77.1 85.1 86.3 93.9 86.0 87.7 95.1 88.0 87.4
-Layer-wise AdaMerging [63] 76.7 87.6 84.8 93.8 85.9 88.1 95.2 88.6 87.6
-We begin by comparing the performance of multi-task model fusion algorithms on various tasks using different models.
-These experiments provide insights into the effectiveness of different fusion methods in improving the performance of multi-task models.
-We evaluate the performance of multi-task model fusion algorithms on image classification tasks using CLIP models, scene understanding tasks using ResNet-50 models, text classification tasks using GPT-2 models, and text-to-text generation tasks using Flan-T5 models.
-  1. 1.
-    Image Classification Tasks with CLIP Models: We utilize the CLIP-ViT-B/32 and CLIP-ViT-L/14 models from the HuggingFace Library [24].
-The results of merging CLIP-ViT-B/32 models on all eight tasks are provided in Table 3 and Figure 4.
-The results of CLIP-ViT-L/14 models are shown in Table 14.
-  2. 2.
-    Scene Understanding Tasks with ResNet-50 Models: We use the NYUv2 dataset and ResNet-50 models for segmentation, depth estimation, and normal estimation tasks.
-In Figure 6, we illustrate the process of merging ResNet-50 models on these tasks, where the backbones are merged, and the heads are copied separately.
-The results of merging ResNet-50 models on these tasks are shown in Table 4.
-  3. 3.
-    Text Classification Tasks with GPT-2 Models: The results of merging GPT-2 models on seven text classification tasks are shown in Table 12.
-  4. 4.
-    Text-to-Text Generation Tasks with Flan-T5 Models: For LoRA fine-tuned Flan-T5-base and Flan-T5-large models, we merge and unload the LoRA adapters before performing multi-task model fusion.
-The results of merging Flan-T5-base and Flan-T5-large models on all eight tasks are shown in Tables 13 and 15, respectively.
-In the above mentioned tables, we compare the performance of different multi-task model fusion algorithms on various tasks.
-The results of pre-trained models, fine-tuned models, and traditional multi-task learning (MTL) are provided as reference methods.
-Appendix C Generalization Experiments
-Table 16: Generalization results on two unseen tasks when merging ViT-B/32 models on six tasks.
-Method Seen Tasks Unseen Tasks
+Weight Averaging 74.6 84.3 84.1 92.8 86.3 87.4 94.8 88.0 86.5
+Task Arithmetic 76.9 85.4 85.3 93.9 85.8 88.1 95.2 87.8 87.3
+Ties-Merging 77.1 85.1 86.3 93.9 86.0 87.7 95.1 88.0 87.4
+Layer-wise AdaMerging 76.7 87.6 84.8 93.8 85.9 88.1 95.2 88.6 87.6
+We extend our evaluation to text-to-text generation tasks using Flan-T5 models fine-tuned with LoRA on eight tasks from the GLUE benchmark (wang2018glue), including both classification and regression tasks.
+Tables 12 and 13 present results for Flan-T5-base and Flan-T5-large models respectively, demonstrating how model fusion performs in the parameter-efficient fine-tuning paradigm.
+The LoRA adapters are merged into the base models before fusion, creating full fine-tuned models that are then subject to various fusion techniques.
+Notably, while Flan-T5-base fusion methods show modest improvements over simple averaging (78.2% vs.
+individual fine-tuned performance of 84.6%), the SMILE model mixing method achieves remarkable performance (84.0%), nearly matching the individual fine-tuned baseline.
+For Flan-T5-large, all fusion methods demonstrate stronger performance, with the best methods (Task Arithmetic, Ties-Merging, Layer-wise AdaMerging) achieving 87.3-87.6% compared to 89.6% for individual models, suggesting that larger models provide more robust representations for effective knowledge combination across tasks.
+This trend aligns with observations from the vision experiments, reinforcing the hypothesis that model scale facilitates more effective parameter-space fusion.
+E.4 Generalization and Robustness Evaluation
+To further assess the generalization and robustness of multi-task model fusion algorithms, we conduct experiments on unseen tasks and corrupted test sets (or out-of-distribution test sets).
+From the experimental results in Section E.4.1 and E.4.2, we have the following key observations: (1) The performance of all multi-task model fusion methods on unseen tasks is generally lower than their performance on seen tasks.
+This is expected, as the models being fused are not explicitly trained on the unseen tasks.
+(2) A negative transfer is observed in Table 15 on the RESISC45 dataset, where the merged models exhibit lower accuracy compared to the pre-trained model.
+The performance of all multi-task model fusion methods on RESISC45 is lower than the pre-trained model, indicating that the knowledge transferred from the seen tasks may not be beneficial or even harmful to this specific unseen task.
+(3) The performance of all methods drops significantly on certain types of corruptions, such as pixelation and impulse noise.
+This highlights the challenge of maintaining robustness under severe distribution shifts and the need for further research in this direction.
+(4) When the test distribution is corrupted, adaptive methods may overfit to certain tasks, leading to a decrease in overall performance.
+This suggests that adaptive methods may need to be further regularized to improve generalization and robustness.
+E.4.1 Generalization Experiments
+Table 14: Generalization results when merging ViT-B/32 models on six tasks (SUN397, Cars, RESISC45, DTD, SVHN, GTSRB) and evaluating on two unseen tasks (MNIST, EuroSAT).
+Method Seen Tasks (ACC) Unseen Tasks (ACC)
+SUN397 Cars RESISC45 DTD SVHN GTSRB Avg. MNIST EuroSAT Avg.
+Pre-trained 63.2 59.9 60.6 43.9 23.5 30.4 46.9 47.6 45.6 46.6
+Fisher Merging 65.5 67.2 78.2 57.6 84.2 75.9 71.4 71.8 49.4 60.6
+RegMean 68.7 70.0 86.5 65.9 93.9 86.7 78.6 82.2 49.3 65.7
+Task Arithmetic 64.3 63.0 73.2 54.9 84.7 79.5 69.9 75.5 42.6 59.1
+Ties-Merging 68.3 65.5 76.9 54.9 75.4 72.0 68.9 73.1 47.3 60.2
+AdaMerging 68.4 71.9 87.9 69.1 92.2 93.8 80.5 77.7 47.3 62.5
+WEMoE 75.4 77.5 94.3 77.0 96.8 98.7 86.6 78.3 44.0 61.1
+Table 15: Generalization results when merging ViT-B/32 models on six tasks (SUN397, Cars, GTSRB, EuroSAT, DTD, MNIST) and evaluating on two unseen tasks (RESISC45, SVHN).
+Method Seen Tasks (ACC) Unseen Tasks (ACC)
 SUN397 Cars GTSRB EuroSAT DTD MNIST Avg. RESISC45 SVHN Avg.
 Pre-trained 63.2 59.9 30.4 45.6 43.9 47.6 48.4 60.6 23.5 40.1
-Fisher Merging [37] 68.1 67.4 67.2 86.4 58.6 81.6 71.5 60.2 42.5 51.3
-RegMean [28] 69.4 70.5 86.9 97.0 67.1 98.3 81.5 50.2 51.5 50.8
-Task Arithmetic [25] 65.2 63.6 76.1 87.1 56.4 94.2 73.8 52.4 45.2 48.8
-Ties-Merging [62] 68.2 65.9 70.0 81.2 56.0 89.0 71.7 60.3 47.3 53.8
-AdaMerging [63] 69.8 72.4 95.5 95.1 70.7 98.1 83.6 48.7 60.7 54.7
-WEMoE [51] 74.3 78.1 98.8 98.7 75.1 99.5 87.4 47.3 51.3 49.3
-For the generalization experiments, we assess the performance of multi-task model fusion algorithms on two unseen tasks after merging ViT-B/32 models trained on six tasks.
-The performance of various multi-task model fusion methods, including Fisher Merging [37], RegMean [28], Task Arithmetic [25], Ties-Merging [62], AdaMerging [63], and WEMoE [51], is compared across both the seen tasks and unseen tasks.
+Fisher Merging 68.1 67.4 67.2 86.4 58.6 81.6 71.5 60.2 42.5 51.3
+RegMean 69.4 70.5 86.9 97.0 67.1 98.3 81.5 50.2 51.5 50.8
+Task Arithmetic 65.2 63.6 76.1 87.1 56.4 94.2 73.8 52.4 45.2 48.8
+Ties-Merging 68.2 65.9 70.0 81.2 56.0 89.0 71.7 60.3 47.3 53.8
+AdaMerging 69.8 72.4 95.5 95.1 70.7 98.1 83.6 48.7 60.7 54.7
+WEMoE 74.3 78.1 98.8 98.7 75.1 99.5 87.4 47.3 51.3 49.3
+For the generalization experiments, we assess the performance of multi-task deep model fusion algorithms on two unseen tasks after merging ViT-B/32 models trained on six tasks.
+The performance of various multi-task model fusion methods, including Fisher Merging (matena2022merging), RegMean (jin2022dataless), Task Arithmetic (ilharco2022editing), Ties-Merging (yadav2023resolving), AdaMerging (yang2023adamerging), and WEMoE (tang2024merging), is compared across both the seen tasks and unseen tasks.
 Specifically, we conduct two sets of generalization experiments using the CLIP-ViT-B/32 models:
@@
 The results are shown in Table 14.
-  • In the second set of experiments, we merge models trained six tasks (SUN397, Cars, GTSRB, EuroSAT, DTD, MNIST) and evaluate the fused model on the unseen tasks (RESISC45, SVHN).
+  • In the second set of experiments, we merge models trained on six tasks (SUN397, Cars, GTSRB, EuroSAT, DTD, MNIST) and evaluate the fused model on the unseen tasks (RESISC45, SVHN).
 The results are shown in Table 15.
-By conducting these two sets of generalization experiments, we aim to gain a comprehensive understanding of how the CLIP-ViT-B/32 models, when fused with knowledge from different task combinations, can perform on various unseen tasks.
-From these experimental results, we can observe instances of negative transfer when evaluating the fused CLIP models on unseen tasks.
-Here, negative transfer occurs when the knowledge gained from fine-tuning on a set of tasks hinders the model’s performance on new, unseen tasks.
-In other words, the model’s ability to generalize and adapt to novel challenges is compromised due to the specific knowledge acquired during the fine-tuning process.
-The presence of negative transfer in these experiments highlights the challenges and limitations of model fusion and generalization in the context of the CLIP-ViT-B/32 models.
-Several factors can contribute to negative transfer, such as:
-  • Task dissimilarity: If the unseen tasks are significantly different from the tasks used for fine-tuning, the learned representations may not be directly applicable, leading to performance degradation.
-  • Overspecialization: Fine-tuning on a specific set of tasks may cause the model to overfit to task-specific features and patterns, reducing its ability to generalize to new tasks.
-  • Interference between tasks: When merging knowledge from multiple tasks, there may be conflicts or interference between the learned representations, hindering the model’s ability to adapt to unseen tasks effectively.
-To mitigate the negative transfer and improve the generalization ability of merged models, several strategies can be explored, such as:
-  • Task selection: Carefully selecting tasks that are more similar or complementary to the target unseen tasks can help reduce the risk of negative transfer.
-This is adapted in Wu et al.
-[58], where the Fisher information matrix is computed for a proxy metric for task similarity.
-  • Regularization techniques: Applying regularization methods, such as weight decay or dropout, during the fine-tuning process may help prevent overfitting and promote better generalization.
-Appendix D Prompt-Based Text-to-Text Generation
-This section details the prompt templates employed for each of the eight text-to-text generation tasks from the GLUE benchmark., see Section 3.3 for more details.
+Tables 14 and 15 present the generalization performance of various multi-task model fusion algorithms when merging CLIP-ViT-B/32 models trained on six seen tasks and evaluating their performance on two unseen tasks.
+This analysis helps us understand how well the fused models can adapt to new tasks that were not encountered during the training and model fusion process.
+E.4.2 Robustness Experiments
+Table 16: Ablations of the test data distribution on ViT-B/32 (for all methods, $\lambda=0.3$).
+Here underlined method names indicate methods that require test-time adaptation training, while italicized method names indicate model mixing approaches.
+Method Cars EuroSAT RESISC45 GTSRB Avg. Cars EuroSAT RESISC45 GTSRB Avg.
+Clean Test Set Corrupted Test Set (Motion Blur)
+Fisher Merging 66.0 92.7 83.7 78.7 80.3 60.7 57.6 81.7 78.4 69.6
+RegMean 72.1 97.5 88.9 93.9 88.1 70.0 71.3 87.5 86.8 78.9
+Task Arithmetic 64.6 91.8 80.2 74.8 77.9 62.4 59.2 78.5 63.3 65.9
+Ties-Merging 65.2 83.3 78.1 67.4 73.5 64.4 53.9 76.4 57.1 62.9
+AdaMerging 75.2 94.3 87.6 96.7 88.5 72.4 72.7 85.3 94.3 81.2
+WEMoE 77.4 98.9 94.4 99.0 92.4 76.5 74.2 93.7 97.4 85.5
+Corrupted Test Set (Impluse Noise) Corrupted Test Set (Gaussian Noise)
+Fisher Merging 61.5 50.0 74.7 52.6 59.7 61.6 48.1 76.0 51.3 59.3
+RegMean 66.9 51.0 80.6 68.7 66.8 69.4 41.8 84.0 67.7 65.7
+Task Arithmetic 59.8 53.3 72.3 45.0 57.6 61.5 52.5 75.0 50.1 59.8
+Ties-Merging 60.2 45.6 69.8 38.3 53.5 61.8 47.3 73.1 42.3 56.1
+AdaMerging 69.2 40.0 79.6 83.3 68.0 70.0 53.3 82.1 80.0 71.4
+WEMoE 75.1 9.7 91.5 91.8 67.0 76.5 9.6 92.7 88.7 66.8
+Corrupted Test Set (Pixelate) Corrupted Test Set (Spatter)
+Fisher Merging 2.2 34.0 17.0 63.2 29.1 61.4 64.2 74.6 47.3 61.9
+RegMean 2.3 38.3 18.2 89.4 37.0 67.7 60.0 81.3 81.9 72.7
+Task Arithmetic 2.3 33.2 19.1 65.6 30.0 61.0 62.5 72.8 57.0 63.3
+Ties-Merging 3.3 31.8 18.0 58.5 27.9 61.3 52.9 70.3 48.1 58.2
+AdaMerging 1.3 52.9 21.0 91.0 41.5 68.4 55.9 78.3 92.3 73.7
+WEMoE 0.5 11.6 2.3 97.5 28.0 75.1 9.7 91.4 96.3 68.1
+Corrupted Test Set (Contrast) Corrupted Test Set (JPEG Compression)
+Fisher Merging 63.8 58.4 75.5 70.4 67.0 66.3 67.6 82.6 58.9 68.8
+RegMean 69.6 64.8 84.4 90.0 77.2 71.5 72.6 88.7 82.2 78.7
+Task Arithmetic 62.3 55.7 75.3 70.8 66.0 63.9 66.1 80.1 61.0 67.8
+Ties-Merging 64.2 52.4 74.8 63.5 63.7 65.0 59.5 77.9 53.2 63.9
+AdaMerging 73.1 67.4 83.0 96.2 79.9 72.9 70.7 86.3 90.6 80.1
+WEMoE 77.2 34.7 93.1 98.4 75.9 77.3 61.0 94.1 95.7 82.0
+To assess the robustness and real-world applicability of multi-task model fusion algorithms, we conduct evaluations on corrupted test sets as presented in Table 16.
+These experiments are crucial for understanding how fusion methods perform under distribution shifts and data quality degradation, which are common challenges in practical deployment scenarios.
+We evaluate six different corruption types across four representative image classification tasks (Cars, EuroSAT, RESISC45, GTSRB), including motion blur, impulse noise, Gaussian noise, pixelation, spatter, and JPEG compression.
+The results reveal varying degrees of robustness across different fusion methods, with some approaches maintaining reasonable performance under mild corruptions while experiencing significant degradation under severe distortions such as pixelation.
+Notably, test-time adaptation-based methods like AdaMerging and WEMoE demonstrate better resilience compared to simpler averaging techniques, though all methods show substantial performance drops under extreme corruption conditions, highlighting the ongoing challenge of maintaining model robustness in multi-task fusion scenarios.
+E.5 Scaling to Large-Scale Neural Networks
+Table 17: Comparison of individual Mistral-7B models and the upscaled model on various benchmark tasks.
+For our method, we set $k_{gate}=8,k=512$, and the total parameter count is 11.2B.
+For a better comparison, we also include the performance of the Qwen1.5-14B model as a reference.
+Model MMLU TruthfulQA GSM8K ARC Challenge
+Mistral-7B-v0.1 (pre-trained) 59.64 42.62 38.81 53.92
+Qwen1.5-14B (reference) 66.11 52.00 69.37 49.93
+MetaMath-Mistral-7B 60.56 44.79 71.49 51.02
+dolphin-2.1-mistral-7b 60.56 55.88 56.93 57.00
+speechless-code-mistral-7b-v1.0 61.18 47.47 48.98 57.68
+Simple Average 61.42 49.95 67.40 57.59
+Task Arithmetic ($\lambda=0.3$) 61.29 49.38 66.94 57.94
+SMILE Upscaled model ($k=512$, 11.2B) 60.66 52.79 67.85 54.35
+SMILE Upscaled model (Dense experts) 60.61 54.23 70.66 55.12
+Model fusion methods can also be applied to large-scale neural networks including Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs).
+The high computational cost associated with developing LLMs are a significant practical challenge for many researchers.
+In FusionBench, we have developed multiple model fusion techniques applicable to LLMs for cheap and efficient model scaling, which can save computational resources for developing new models.
+Moreover, FusionBench is actively developing and integrating more LLM and MLLM merging methods, with ongoing work to expand benchmark coverage to large-scale model fusion tasks.
+While evaluation methodologies for LLMs and MLLMs are still maturing and computational costs remain a challenge, FusionBench supports established evaluation frameworks such as LM-Evaluation-Harness (eval-harness) for assessing model performance.
+As the toolkit evolves, more comprehensive LLM task evaluations will be included in future releases.
+Take merging Mistral-7B models using SMILE upscaling (tang2024smile) as an example, we compare the performance of the individual Mistral-7B models and the upscaled model on various benchmark tasks in Table 17, as well as the performance of the Qwen1.5-14B model as a reference.
+The LM-Evaluation-Harness is utilized to assess the performance of the models.
+We fuse three Mistral-7B models, each fine-tuned for a distinct downstream task as showing varying performance across different tasks, thereby incorporating task-specific expertise.
+It is observed that the upscaled models and merged methods (Simple Average, Task Arithmetic) generally enhance performance compared to individual models, demonstrating the benefits of model fusion techniques.
+Appendix F Prompt-Based Text-to-Text Generation
+This section details the prompt templates employed for each of the eight text-to-text generation tasks from the GLUE benchmark, see Section E.3 for more details.
 Within each task, we provide the format of the input text, and the corresponding target text mapping.
@@
     • Target Text: ":.1f", parse to float with one decimal place
+Reporting Metrics:
+We report accuracy for all tasks except for STSB, where we use Spearman’s $\rho$ as the evaluation metric.
+For task STSB, the model is expected to output a numerical value.
+An example from the STSB task is as follows:
+  • Input:
+    • Sentence 1: A plane is taking off.
+    • Sentence 2: An air plane is taking off.
+  • Output:
+    • label: 5
+We try to parse the output as a numerical value.
+If the model outputs a numerical value, we can calculate the Spearman’s rho between the predicted numerical value and the ground truth numerical value.
+If the model outputs a non-numerical value, we assume the Spearman’s rho is 0, indicating that there is no discernible monotonic increasing or decreasing relationship between the model’s predictions and the ground truth.
+This is a conservative approach, as even non-numerical outputs might contain some relevant information that’s being discarded in this evaluation.